### v1

# Nash CoT: Multi-Path Inference with Preference Equilibrium

Abstract
Chain-of-thought (CoT) prompting has emerged as a powerful technique for enhancing the reasoning capabilities of Large Language Models (LLMs) on complex problems. Among CoT-related studies, self-consistency (Multi-path inference with answer filtering through voting) involves generating multiple reasoning paths using the CoT framework and then selecting the most frequently produced outputs standing out as a concise yet competitive approach. While self-consistency has indeed led to the improvements in LLM inference, the use of multi-path inference also escalates deployment costs. Therefore, maintaining the performance benefits of self-consistency inherited from multi-path inference while reducing the inference costs holds significant value. In this research, we conceptualize language decoding as a preference consensus game, constructing a bi-player gaming system within each local path, and introduce Nash Chain-of-Thought (Nash CoT). Specifically, for a given question, we leverage LLM to autonomously select the contextually relevant template and generate outputs guided by this template, aiming to reach Nash Equilibrium alongside normal generation in each path. This approach allows us to achieve comparable or improved performance compared to self-consistency while using fewer inference paths on various inference tasks, including Arabic reasoning, Commonsense Question answering, and Symbolic inference.
Nash CoT: Multi-Path Inference with Preference Equilibrium
Ziqi Zhang1thanks:    Equal contribution.,  Cunxiang Wang [^1] footnotemark: 1,  Xiong Xiao2,  Yue Zhang1thanks:    Corresponding authors.,  Donglin Wang122footnotemark: 2 1 School of Engineering, Westlake University 2 University of Cambridge {zhangziqi, wangcunxiang, zhangyue, wangdonglin}@westlake.edu.cn
1 Introduction
Large Language Models (LLMs) have revolutionized Natural Language Processing (NLP) (Ouyang et al., 2022; etc., 2023; Jiang et al., 2023; Brown et al., 2020b; OpenAI, 2024). In particular, leveraging human-designed instructions as input, LLM demonstrates superior inference performance across various types of simple reasoning tasks (Radford et al., 2019; Brown et al., 2020a). But, its performance remains variable in complex reasoning tasks (Rae et al., 2022). To enhance LLM’s inference capabilities on complex issues, we can employ a step-by-step inference approach, known as Chain-of-Thought (CoT) prompting (Wei et al., 2023). For instance, starting with a template like "Let’s think step by step", LLM first generates rationales and then arrives at a final prediction. This approach significantly improves LLM’s inference performance across tasks like Arabic, Symbol Inference, and CommonsenseQA.
Figure 1: Demonstrations of Self-Consistency and Nash Chain-of-Thought (Nash CoT). (a) Self-Consistency entails inferring multiple paths and subsequently voting the most frequently prediction. (b) In Nash CoT, inference occurs multiple times, with each path generating two responses, but only one reached Preference Equilibrium is sustained. Ultimately, the answer sampled is the one that reaches preference equilibrium.
Subsequently, the impressive performance of CoT on complex inference tasks has spurred new developments in this direction (Wang et al., 2023; Wei et al., 2023; Jin and Lu, 2023; Zhang et al., 2022; Shi et al., 2022). Among these developments, self-consistency (Wang et al., 2023) emerges as a competitive CoT approach. It leverages uncertainty from multiple inference paths from a single LLM, and ranking generated answers by frequency can significantly enhance LLM’s inference performance. This approach significantly improves the performance of GPT-3 utilizing zero-shot CoT without any options for parameter tuning. Meanwhile, experimental results indicate that inference performance improves with the number of generated samples (Wang et al., 2023), suggesting that the potential of LLM’s inference capabilities has not yet to be fully exploited. Although self-consistency is straightforward to implement and requires no additional turning, it has a significant drawback: substantially higher inference costs compared to directly utilizing CoT.
On the other hand, the performance improvements resulting from self-consistency imply that relying solely on single-path inference cannot fully harness the inference capabilities of LLM, and multi-path inference encompasses the potential correct answers. Therefore, it’s necessary to maintain this strategy. Meanwhile, numerous practical applications indicate that when a LLM is given the appropriate templates, it can perform specific tasks more proficiently. Hence, an intuitive strategy to reduce the number of inference paths in self-consistency is to use templates to guide the LLM to correctly infer each path.
Figure 2: General Performance Comparison. We compare the average performance of , zero-shot, and zero-shot CoT self-consistency (20 Paths) with our Nash CoT (10 Paths) on Mistral-Instruct and GLM4.
To achieve this goal, we use the contextual information required by the question as a template to guide the inference in each path. This approach increases the probability that the LLM can correctly solve the question in each path, thereby potentially reducing the number of inference paths needed for multi-path inference. Meanwhile, to alleviate overconfidence in template-guided generation, we develop a bi-player gaming system that introduces the Nash Equilibrium of preference of strategy (defined as Preference Equilibrium) in multi-path inference. This system samples generations that balance the preferences of both template-guided and normal generation, ensuring the output robustly matches the context of the given question while sustaining some moderate generation pattern. Subsequently, we combine multi-path inference with Preference Equilibrium to propose Nash CoT. We present the comparison between Self-Consistency and Nash CoT in Figure 1.
We conduct experiments with two local deployed LLMs-Mistral-Instruct (Jiang et al., 2023) and GLM4 (Zeng et al., 2022; Du et al., 2022) on various inference tasks, including Arabic Reasoning (Koncel-Kedziorski et al., 2015; Hosseini et al., 2014) and Symbolic Reasoning (Wei et al., 2023) and Commonsense Reasoning (Talmor et al., 2019; Geva et al., 2021). As shown in Figure 2, Nash CoT has achieved similar or even better performance against self-consistency with fewer inference paths. Meanwhile, as shown in Figure 3, Nash CoT has significantly reduced the inference cost by up to 50% on local deployed LLMs.
To our knowledge, we are the first to introduce Preference Equilibrium into multi-path inference with the aim of harmonizing text generation guided by the optimal template with that generated by the model’s default state. It can reduce the number of paths needed to achieve good results in multi-path inference. And we are also the first to integrate Preference Equilibrium into multi-path inference and propose Nash CoT. It can achieve similar or even superior performance on various reasoning tasks while requiring only half costs compared to self-consistency. [^1] We will release our code at ¡https://github.com/stevezhangzA/nash-chain-of-thought¿.
2 Related Work
Chain-of-Thought (CoT).
There are three majority kinds of CoT approaches. Zero-shot CoT that Prompts LLM with simple yet instructions to guide LLM generate step by step (Kojima et al., 2023). Manual CoT that initialized by randomly sampled several cases from dataset or designed by human, and followed by utilizing these cases as demonstration to guided LLM generate follow the manner of demonstration (Wei et al., 2023), however, such methods can be biased if the demonstration can represent real distribution. Automatic (or batch) CoT (Zhang et al., 2022) first sample cases which have the most representative sentence embedding in each clusters, followed by inference with the same manner as manual CoT. Self-Consistency (Wang et al., 2023) showcases strong performance in vast benchmarks. Apart from its impact on inference performance,self-consistency also boasts scalability as a key advantage. It seamlessly integrates with different approaches, such as tabular consistency of transformations (tab-CoT) Jin and Lu (2023), making it adaptable and versatile for various applications. Despite self-consistency has improved LLM’s performance on Arabic benchmarks. Correspondingly,self-consistency should have to inference multi-times, thus it burdens the deploy budgets.
One way to address this limitation is by initially sampling multiple inference paths and then fine-tuning using the highest frequency path. Specifically, Huang et al. (2022) suggest that gathering inferences from multiple paths and sampling the most frequent generation to fine-tune a smaller LLM can enhance the LLM’s inference performance. However, this approach still requires updating the LLM’s parameters, which is inefficient. Therefore, it is necessary to further explore inference methods to maintain the performance of self-consistency while reducing the number of inference paths.
Preference Optimization.
The training policy with Reinforcement Learning (RL) to reflect preference, termed Reinforcement Learning with Human Feedback (RLFH), was initially introduced by Akrour et al. (2011) and has since undergone consistent improvement and refinement by Cheng et al. (2011); Busa-Fekete et al. (2013); Wilson et al. (2012). This approach has been widely applied to adjust Large Language Models’ (LLMs) parameters to align with human preferences (Ouyang et al., 2022; Jiang et al., 2023; etc., 2023). Recently, a new approach called Direct Optimizing from Preference (DPO) has been proposed by Rafailov et al. (2023), aiming to directly adjust LLMs to reflect human preferences without requiring a reward model and RL algorithm. Additionally, Munos et al. (2023) proposed combining DPO with Nash equilibrium to ensure convergence of the last iterated policy. Our study also utilizes the concept of equilibrium in preference model, but the main difference is that we utilize preference equilibrium as a standard to pick up the most preferred answer instead optimizing the LLM’s parameters.
3 Preference Equilibrium in mini-batch inference
Self-consistency has shown that the inference of LLMs under a single path may not represent their full capabilities. By simply conducting multiple inferences and filtering answers through voting, it is possible to achieve more accurate results. However, multi-path inference lacks a strong theoretical foundation to determine the optimal number of inference paths, potentially leading to much more computational resource consumption. To reduce the number of paths required for multi-path inference, we utilize the concept of Nash Equilibrium to locally construct a binary game system in multi-path inference. Specifically, the preference of each valid inference path of the LLM needs to achieve Nash equilibrium with the preferences of the generation guided by the template. This approach increases the probability of each path correctly answering the question while maintaining a certain level of robustness, thereby reducing the number of inference paths required by self-consistency.
Preference Model.
Given text input $x$, and the sampled predictions (answers) $y_{1}$, $y_{2}$, we first define $y_{1}$ prefers $y_{2}$ as Equation 1:
$\begin{split}\mathcal{P}(y_{1}\prec y_{2}|x):={\rm sign}(r_{\theta}(y_{1}|x))-{\rm sign}(r_{\theta}(y_{2}|x)),\end{split}$ (1)
[…]
Player Templates: Templates for our preference model that are shown the structure: id(player): description. Mathematician: You are a mathematician, you excel at analyzing problems from a mathematical logical perspective and arrive at conclusions that align with your values. Literary scholar: You are a literary scholar who has read a vast array of literary works. Please consider the problem from the perspective of a literary scholar.
Philosophical: You are a philosopher, your knowledge base includes a wealth of philosophical knowledge. You enjoy approaching problems from a philosophical perspective and arriving at conclusions that align with your values.
Geographer: You are a geographer with a deep understanding of geographical knowledge. Please approach the given problem from the perspective of a geographer.
$\cdots$ (other cases have been appended to the Appendix.)
Additionally, drawing from the definition from Munos et al., we define one policy as more preferred over another as:
$\begin{split}\mathcal{P}(\pi_{1}\prec\pi_{2}):=\mathbb{E}_{{}^{y_{1}\sim\pi_{1}(\cdot|x)}_{y_{2}\sim\pi_{2}(\cdot|x)}}[\mathcal{P}(y_{1}\prec y_{2}|x)]\end{split}$ (2)
Preference Equilibrium.
We aim to select the best template (we provide several cases in above example 3 for the current problem, thus facilitating the large model in problem-solving, correspondingly reduces the required num of inference paths. However, this may lead to some issues: 1) If the template is incorrectly chosen, it may cause the agent to generate answers outside the range of correct responses corresponding to the current problem, resulting in erroneous replies. 2) The large model may excessively generate context-dependent responses, affecting its robustness. To address these issues, we build a local bi-player gaming system that the preference of template guided LLM (player 1) over normal status of LLM (player 2) is the pay-off of template guided LLM, vice visa. If player 1 and player 2 reaches Nash Equilibrium  [^2] Preference Equilibrium leverages the concepts of Nash Equilibrium. Nash equilibrium is proposed by John Nash. It is a concept solution where, assuming each participant knows the equilibrium strategies of the other participants, no participant can benefit by changing their own strategy., then the strategy can match the preference of both player 1 and player 2.
Subsequently, we define the status that player 1 and player 2 reach Nash Equilibrium as Preference Equilibrium (Definition 1). Meanwhile, in Theorem 3.1, we prove the existence of Nash Equilibrium in this system. Specifically, the strategy of player 1 equal to player 2 is one solution that this system has reached Preference Equilibrium.
Theorem 3.1 (Existence of Preference Equilibrium).
Given any two policy (player) $\pi_{1}$ and $\pi_{2}$ within the gaming system defined in Definition 1, where $\pi\in\Pi$. $\pi_{1}\equiv\pi_{2}$denotes a solution where the gaming system reaches Nash Equilibrium.
Proof of Theorem 3.1 see Appendix E.
Meanings of the existence of Preference Equilibrium.
Theorem 3.1 proves the existence of an Nash Equilibrium between the template guided LLM and the normal status of LLM. When reaching Preference Equilibrium, the preference of decisions made by the template guided LLM are aligned with those made by the LLM under normal status. Meanwhile, the preference of template guided LLM generation is much more closed to the requirement of quesitons’ context, while those of the normal status LLM are predominantly based on its parameters which is much more robust than template guided LLM. Therefore, this equilibrium can balance the requirement of contextual information and robustness of the model generation during problem-solving. Notabaly, $\pi_{1}\equiv\pi_{2}$ means their outputs are also likely to be equal. This insight forms a fundamental basis for piratically implementing Nash CoT.
3.1 Mini-batch inference with Preference Equilibrium
Subsequently, based on the concept of Preference Equilibrium, we conceptualize a Mini-batch inference (shown in Figure 1) as a bi-player gaming system. This approach aims to achieve better inference compared to direct inference, while still preserving some of the inherent randomness of standard inference methods. To begin with proposing this system, we first define $x^{t}$ as the template of zero-shot CoT, $\{x^{c}_{0},x^{c}_{1},\cdots,x^{c}_{n}\}$ (we provided several cases in Player Templates) as the candidates template for template guided generation. Meanwhile, in this system, the template can to be chosen by a reward model $r_{\theta}$.
[…]
4 Nash Chain of Thought (Nash CoT)
Algorithm 1  Nash CoT (Answer Gathering)
Require: Candidate question q sampled from $\mathcal{Q}=\{q_{0},q_{2},\cdots,q_{n}\}$; Outer iterations ${\rm n}_{\rm outer}$; Num of mini-batch inference $\rm n_{mini}$; Large language model $\pi$; CoT prompt $x^{t}$, candidiate player template $\{x^{c}_{0},x^{c}_{1},\cdots,x^{c}_{n}\}$, ${\rm Prompt}(\{x^{c}\})$ is used to point out the most preferred $x^{c}$.
Generation:
1:  Initialize answer list ${\rm ans}=[\,]$.
2:  $x_{c}\leftarrow\pi(\cdot|{\rm Prompt}(\{x^{c}\}))$
3:  for t in range(${\rm n}_{\rm outer}$do
4:     Initialize prefer pairs ${\rm pref}=[\,]$.
5:     for t in range($\rm n_{mini}$do
6:        $y\leftarrow\pi(\cdot|x^{t},q)$;
7:        ${\rm ans}.$append($y$);
8:     end for
9:     $y^{*}\leftarrow\pi(\cdot|x^{c},x^{t},x)$
10:     $\tau.{\rm append}([y^{*},{\rm ans}])$
11:  end for
12:  Return $\tau$
Nash CoT can be seen as an extension of Mini-batch inference with Preference Equilibrium, implementing multiple Mini-batch inferences to enhance performance. This approach is influenced by experimental results from self-consistency, which suggest that increasing the number of paths can improve inference accuracy. Meanwhile, the reasoning process for each question is divided into two stages: Answer Gathering and Answer Filtering.
Answer Gathering.
When generating candidate answers, the process predominantly involves two types of loops: Mini-batch Loops ($\rm n_{\rm mini}$): In Chapter 3.1, we discussed the implementation of mini-batch inference with Preference Equilibrium. As shown in Algorithm 1, this process involves searching for template-guided generations within two rounds of generation ($[y_{1},y_{2}]$). We refer to the times of these two predictions as the $\rm n_{\rm mini}$. Moreover, to mitigate the impact brought from low-frequency predictions, we introduce iterating $\rm n_{\rm mini}$ multiple times. This leads us to another type of loop: Outer Loops ($\rm n_{\rm outer}$): This loop resembles the concept of multi-path in self-consistency. After completing loop $\rm n_{\rm outer}$, we filter the generated answers and retain the answer that reaches equilibrium most frequently (shown in Algorithm 2), as the preferred answer.
Core LLM Methods SingleEQ AddSub MultiArith GSM8K AQuA SVAMP Avg.
Mistral-Instruct (7B) zero-shot 15.3 $\pm$ 0.8 12.0 $\pm$ 2.8 3.3 $\pm$ 1.3 2.7 $\pm$ 2.0 20.8 $\pm$ 1.5 7.7 $\pm$ 2.0 10.3$\pm$ 6.5
zero-shot CoT 76.0 $\pm$ 0.8 82.5 $\pm$ 2.0 75.4 $\pm$ 6.1 44.3 $\pm$ 4.0 27.9 $\pm$ 2.3 63.4 $\pm$ 6.9 61.6$\pm$19
self-consistency (20 Paths) 82.5 $\pm$ 0.8 86.3 $\pm$ 5.1 86.3 $\pm$ 2.8 58.5 $\pm$ 2.8 34.4 $\pm$ 6.1 76.5 $\pm$ 2.8 70.8$\pm$ 19
Nash CoT (10 Paths) 81.4 $\pm$ 0.8 86.3 $\pm$ 6.0 86.3 $\pm$ 4.7 55.7 $\pm$ 5.8 39.9 $\pm$ 5.4 77.0 $\pm$ 3.5 71.1$\pm$ 17
GLM4-chat (9B) zero-shot 1.1 $\pm$ 1.5 1.1 $\pm$ 1.5 12.6 $\pm$ 3.9 12.0 $\pm$ 2.0 22.4 $\pm$ 4.1 4.4 $\pm$ 2.8 8.9$\pm$7.6
zero-shot CoT 90.7 $\pm$ 1.5 90.7 $\pm$ 1.5 98.4 $\pm$ 1.3 80.9 $\pm$ 2.8 20.8 $\pm$ 3.1 86.9 $\pm$ 3.5 78.1$\pm$26.1
self-consistency (20 Paths) 92.3 $\pm$ 2.0 90.2 $\pm$ 2.3 98.4 $\pm$ 2.3 89.3 $\pm$ 0.2 20.8 $\pm$ 3.1 91.5 $\pm$ 1.1 80.4$\pm$26.8
Nash CoT (10 Paths) 91.3 $\pm$ 0.8 90.2 $\pm$ 2.7 96.7 $\pm$3.3 80.3$\pm$ 1.3 20.8$\pm$3.1 88.0 $\pm$2.0 77.9$\pm$26.1
Table 1: Experimental results on arithmetic reasoning benchmarks. We test Zero-Shot CoT and Nash CoT with the core LLM includes Mistral-Instruct (7B) and GLM4-chat (9B) on mathematical benchmarks including AddSub, MultiArith, SingleEQ, SVAMP, GSM8K, and AQuA. Nash CoT performs the best.
Answer Filtering.
In terms of answer filtering, as shown in Algorithm 2 we first count the most frequent prediction satisfy Preference equilibrium. Specifically, we count all $y^{*}$ satisfy $y^{*}\in[y_{1},y_{2}]$ and compute their frequency. Subsequently, we return the most frequent case. Otherwise, if is no cases satisfy $y^{*}\in[y_{1},y_{2}]$, we adopt the strategy of self-consistency by selecting the most frequent prediction among all generated answers.
Algorithm 2  Nash CoT (Answer Filtering)
Require: Preference pair list $\tau$
Filtering:
1:  Initialize hash table: ${\rm hash}=\{\}:k\rightarrow v$.
2:  Initialize new answer list ${\rm n\-ans}=[\,]$.
3:  for $[y^{*},{\rm ans}]_{i}$ in $\tau$ do
4:     if $y^{*}\in{\rm ans}$ then
5:        ${\rm hash}$[$y^{*}$]+=1
[…]
7:  end for
8:  if ${\rm hash}\equiv\{\}$ then
9:     return  the most frequent y in ${\rm n\-ans}$
10:  else
11:     return  $y\leftarrow k=\arg\max_{v}{\rm hash}$
12:  end if
Subsquently, we propose Nash CoT, which iterates through Algorithm 1 and Algorithm 2 to perform inference on all sampled questions, where $\tau$ represents the candidate answers from the Answer Gathering stage.
[…]
Practical Implementation of reward model $r_{\theta}$.
In the process of practical implementation, we do not explicitly train a reward model $r_{\theta}$ to confine the player template $x^{c}$ (we have provided cases in Player Templates) using Equation 1. Instead, we directly use the preference template (shown in Preference Template) to guide the LLM in determining the most suitable player template for a given question. For example, when presented with a coin flip question as shown in Figure 4, we fill the Preference Template with given question and player templates. This filled template is then input into the LLM to provide the id of the most suitable player template from the available options. In particular, we believe it’s effective, this is because most of baselines we selected have been turned to reflect human preference, thus we believe the selected LLM can be directly utilized as the preference model to point out the most preferred option among candidate options.
5 Experiments
Core LLM Methods Coin-Flipping Last Letters Object Tracking Bigbench Date Avg.
Mistral-Instruct (7B) zero-shot 26.8 $\pm$ 5.1 0.0 $\pm$ 0.0 35.5 $\pm$ 4.1 31.1 $\pm$ 7.6 23.4$\pm$ 14
zero-shot CoT 27.9 $\pm$ 4.0 0.0 $\pm$ 0.0 30.1 $\pm$ 2.8 36.6 $\pm$ 5.4 23.6$\pm$ 14
self-consistency (20 Paths) 21.9$\pm$ 4.7 0.0$\pm$ 0.0 38.8$\pm$ 0.8 47.0$\pm$ 1.5 26.9$\pm$ 18
Nash CoT (10 Paths) 29.0$\pm$ 5.4 0.5 $\pm$ 0.8 44.8 $\pm$ 2.0 41.1 $\pm$ 1.2 28.9$\pm$ 17
GLM4-chat (9B) zero-shot 27.3 $\pm$ 6.3 0.0 $\pm$ 0.0 38.8 $\pm$ 0.8 16.4 $\pm$ 2.3 20.7$\pm$14.3
zero-shot CoT 87.4 $\pm$ 0.8 0.0 $\pm$ 0.0 37.7 $\pm$ 2.3 16.4 $\pm$ 4.8 35.4$\pm$32.9
self-consistency (20 Paths) 98.9$\pm$1.5 0.0$\pm$0.0 37.7$\pm$2.3 16.4$\pm$4.8 38.3$\pm$37.5
Nash CoT (10 Paths) 93.4$\pm$2.7 0.0 $\pm$ 0.0 37.7$\pm$2.3 16.4$\pm$4.8 36.9$\pm$35.3
Table 2: Experimental results on symbolic inference benchmarks. We test Zero-Shot CoT and Nash CoT with Mistral-Instruct (7B) and GLM4-chat (9B) on Symbolic QA benchmarks includes Coin-Flipping, Last Letters and Object Tracking. Among these baselines, Nash CoT performs the best.
Core LLM Methods StrategyQA CommonsensQA Avg.
Mistral-Instruct (7B) zero-shot 49.2 $\pm$ 8.8 62.3 $\pm$ 4.8 55.8$\pm$ 7
zero-shot CoT 57.4$\pm$2.3 70.5 $\pm$ 2.7 64.0$\pm$7
self-consistency (20 Paths) 59.6 $\pm$ 2.0 71.0 $\pm$ 3.4 65.3$\pm$ 6
Nash CoT (10 Paths) 56.8 $\pm$ 2.0 69.4 $\pm$ 4.7 63.1$\pm$ 6
GLM4-chat (9B) zero-shot 56.8 $\pm$ 4.7 17.5$\pm$ 2.0 37.2$\pm$19.7
zero-shot CoT 63.9$\pm$2.3 18.0$\pm$2.3 41.0$\pm$ 22.9
self-consistency (20 Paths) 69.9 $\pm$3.3 18.0$\pm$2.3 44.0$\pm$26.0
Nash CoT (10 Paths) 66.7 $\pm$ 0.8 18.0 $\pm$ 2.3 42.4$\pm$24.4
Table 3: Experimental results on Commonsense Reasoning. We test Zero-Shot CoT and Nash CoT with Mistral-Instruct (7B) and GLM4-chat (9B) on Commonsense Reasoning datasets includes StrategyQA and CommonsenseQA .
The goal of our experiment is to 1) demonstrate the performance advantage and effectiveness of Nash CoT. 2) shocase whether Nash CoT help reduce the overall inference time. In the following sections, we first introduce our experimental setup and then present the experimental results and analysis.
Datasets.
Our majority benchmarks are composed of three different kinds of inference tasks. 1) arithmetic reasoning: SingleEq (Koncel-Kedziorski et al., 2015), AddSub (Hosseini et al., 2014), MultiArith (Roy and Roth, 2016), GSM8K (Cobbe et al., 2021), AQUA (Ling et al., 2017), and SVAMP (Patel et al., 2021). 2) symbolic reasoning: Last Letters, Coin Flip (Wei et al., 2023), and Object Tracking, Bigbench Date. 3) commonsense question answering: CommonsenseQA Talmor et al. (2019) and StrategyQA Geva et al. (2021). For more details about the dataset please refer to Appendix A.
LLMs.
To validate that Nash CoT is a general CoT method, we selected different large models as test models, including Mistral-7B (Instruct) (Jiang et al., 2023), GLM4-9B-Chat Zeng et al. (2022); Du et al. (2022). In particular, all of these selected LLMs are turned via RL with human feedback (RLHF), and the difference between LLM turned with RLHF and the original foundation models have been detailed by Ouyang et al. (2022).
Baselines.
The preliminary baselines we utilized include zero-shot, zero-shot CoT (Wei et al., 2023), andself-consistency (Wang et al., 2023). We test these approach with freezed LLMs.
Settings.
Our evaluation on all selected tasks utilizes the same experimental settings bellow:
  • zero-shot and zero-shot CoT. We follow the method proposed by Wei et al. (2023) and use the original template (e.g., "Let’s think step by step") for evaluation.
  • self-consistency. We follow Wang et al. (2023) to evaluate the performance ofself-consistency with selected LLMs, utilizing the zero-shot CoT template. Additionally, we set the number of inference paths to 20.
  • Nash CoT. We set up $n_{\rm outer}$ as 3 and $n_{\rm mini}$ as 2, resulting in a total of $n_{\rm outer}\times(n_{\rm mini}+1)+1=10$ paths. Additionally, we have provided the player templates $x^{t}$ in Table 1 and the Appendix, meanwhile we utilizing the same CoT template $x^{c}$ as in zero-shot CoT.
Additionally, all evaluations are conducted on the inference of 60 random sampled questions multi times. And we have provided the mean and standard error in all tables.
Figure 3: We used GLM4-chat (9B) on the same type of GPU (A-100) to evaluate Nash CoT and self-consistency across selected tasks. Nash CoT, employing a total of 10 paths, requires nearly half the time of self-consistency, which has 20 paths in total.
5.1 Experimental Results
Evaluated Scores.
The majority experimental results are demonstrated in table 12 and 3. Nash CoT can improve Mistral-Instruct (7B) on almost all selected inference tasks, while showcasing similar performance to self-consistency with twice inference paths on GLM4-chat (9B). In particular, we have provided the total paths of Nash CoT that it only require the half of self-consistency, thus our claim in section 3 can be validated.
Figure 4: We use Mistral-Instruct (7B) to examine the impact of loop numbers on the inference performance of the large language model. Specifically, we used solid lines of specific colors to represent the experimental performance under certain $N_{\rm outer}$ as the $N_{\rm mini}$ changed. We marked self-consistency with 20 paths using dashed lines, and some results of Nash CoT, with total paths close to 20, were marked with stars.
When focusing on Mistral-Instruct (7B), Nash CoT has better performance on arithmetic and symbol inference tasks, showcasing its superiority performance on logic/math inference tasks. However, Nash CoT does not showcase improved performance in commonsense question answering tasks. We argue that this is because commonsense question answering tasks are more diverse, and the player template can’t cover all topics. Therefore, the player template limits Nash CoT on commonsense question answering tasks. Importantly, we limit Nash CoT’s performance by utilizing only total 10 paths for inference in this section. However, additional experimental results in the ablation section show that Nash CoT outperforms self-consistency via increasing the inference loops, thus Nash CoT can outperform self-consistency.
Inference Time.
[…]
6 Ablation Study
In order to further validate the effectiveness of Nash CoT, we conducted extensive ablations to answer the following questions: 1) What will happen when the number of inference paths for Nash CoT is further increased? Will Nash CoT eventually surpass self-consistency, and what is the relationship between the number of loops and performance? 2) Does the template really improve the accuracy of path predictions, and what impact does it have on experimental performance?
As the number of inference paths increases, Nash CoT can obviously surpass self-consistency with fewer inference paths.
To address question 1), we selected Mistral-Instruct (7B) and conducted evaluation on three different reasoning tasks, adjusting the $N_{\rm mini}$ and $N_{\rm outer}$. As shown in Figure 4, as the number of loops increases, Nash CoT has a high probability of significantly outperforming self-consistency with fewer paths. However, different from self-consistency, the experimental results of Nash CoT do not show a monotonic (linear) relationship with the total number of total paths. This indicates that there is a significant difference between Nash CoT and self-consistency. Unlike Nash CoT, the experimental results of self-consistency show a clear improvement in performance as the number of paths increases.
The performance is impacted by the player template.
To illustrate the impact of the template, we removed the mathematical templates from the Player Templates and then evaluated Nash CoT on selected Arabic reasoning. Results are shown in Table 4, showing an approximately 9.2% decrement in GSM8K and 6.2% decrement in SVAMP. Therefore, the performance of Nash CoT is impacted by the Player Template.
GSM8K AQua SVAMP
55.7$\rightarrow$ 50.6 39.9$\rightarrow$ 39.8 77.0 $\rightarrow$72.2
Table 4: Performance decreasing. We remove the mathematics from Player Templates and test Nash CoT on selected Arabic reasoning tasks.
7 Conclusion
In this study, we proved the existence of Nash equilibrium in preference model, subsequently, we proposed a new CoT approach Nash CoT, and validated its performance on various inference benchmarks. Experimental results show that Nash CoT can perform equally or even better than self-consistency while only require half inference costs.
Limitations and Future Work
Despite Nash CoT showcase competitive performance with only half of inference paths, it requires pre-defined template, thus it’s in-convenient to utilize Nash CoT in new emerging scenario, in the future we will develop a automatic approach to balance task feedback and template design.
Ethics Claims
Despite LLM has showcased superiority performance on vast benchmarks, but pre-train or fine-tune a LLM requires numerous computing resources. Therefore, it’s crucial to study how to inference a LLM to reach the ceiling of its capacity. CoT is a ideal approach which has been proved that can obviously evaluate the performance of LLMs’ inference. Among that,self-consistency is one of the best CoT approach.
Our method effectively reduce the inference times of multi-path inference, thereby reducing the deploy budgets of self-consistency. We believe our approach can further elevate the effectiveness of multi-path inference, thereby further improving the effectiveness of LLM.
acl
Appendix A Dataset
Our majority dataset are composed of three different kinds of inference tasks. 1) arithmetic reasoning: SingleEq (Koncel-Kedziorski et al., 2015), AddSub (Hosseini et al., 2014), MultiArith (Roy and Roth, 2016), GSM8K (Cobbe et al., 2021), AQUA (Ling et al., 2017), and SVAMP (Patel et al., 2021). 2) symbolic reasoning: Last Letters, Coin Flip (Wei et al., 2023), and Object Tracking, Bigbench Date. 3) commonsense question answering: CommonsenseQA Talmor et al. (2019) and StrategyQA Geva et al. (2021). For more details about the dataset please refer to Wang et al. (2023).
Appendix B Uesage of LLM.
We utilize LLM to rectify grammar errors.
Appendix C Computing Resources
Our experiments were run on a computer cluster with 32GB RAM, 4-Core CPU, and NVIDIA-A100 (80G, 32G)/NVIDIA-V100 (32G) GPU, Linux platform.
Appendix D Source Code.
We have provided source code for reference. Additionally, our code are based on https://github.com/amazon-science/auto-cot and refer to the coding manner from https://github.com/eureka-research/Eureka.
Appendix E Proof of theorem 3.1.
Subsequently, we prove the existence of Nash equilibrium in this system. For any two given polices $\pi_{1}\in\Pi$ and $\pi_{2}\in\Pi$ We first define the pay-off of $\pi_{1}$ and $\pi_{2}$ as $R(\pi_{1};\pi_{2})$ and $R(\pi_{2};\pi_{1})$:
$\begin{split}R(\pi_{1};\pi_{2})=\mathcal{P}(\pi_{1}\prec\pi_{2})\\ R(\pi_{2};\pi_{1})=\mathcal{P}(\pi_{1}\succ\pi_{2}),\end{split}$ (3)
we provide the proof of the existence of Nash equilibrium in this system. We define $\bar{\pi}=[\pi_{1},\pi_{2}]$, $v(\bar{\pi})=[R(\pi_{1};\pi_{2}),R(\pi_{1};\pi_{2})]$. According to the Nash equilibrium, it should have to satisfy this relationship:
$\begin{split}v(\bar{\pi}^{*})(\bar{\pi}^{*}-\bar{\pi})\leq 0\end{split}$ (4)
Subsequently, refer to Munos et al., we can learn that if we want Equation 4 holds true, we just have to guarantee Equation 5 holds true.
$\begin{split}(v(\bar{\pi})-v(\bar{\pi}^{\prime}))^{\rm T}(\bar{\pi}-\bar{\pi}^{\prime})\leq 0,\end{split}$ (5)
where $\bar{\pi}$ and $\bar{\pi}^{\prime}$ are any two given policy set. Subsequently, we can further darrive at the following relationships:
$\begin{split}(v(\bar{\pi})-v(\bar{\pi}^{\prime}))^{\rm T}(\bar{\pi}-\bar{\pi}^{\prime})&=\begin{pmatrix}R(\pi_{1};\pi_{2})-R(\pi_{1}^{\prime};\pi_{2}^{\prime})\\ R(\pi_{2};\pi_{1})-R(\pi_{2}^{\prime};\pi_{1}^{\prime})\\ \end{pmatrix}\cdot\begin{pmatrix}\pi_{1}-\pi_{1}^{\prime},\pi_{2}-\pi_{2}^{\prime}\\ \end{pmatrix}\\ &=\bigg{(}R(\pi_{1};\pi_{2})-R(\pi_{1}^{\prime};\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime})+\bigg{(}R(\pi_{2};\pi_{1})-R(\pi_{2}^{\prime};\pi_{1}^{\prime})\bigg{)}\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime})+\bigg{\{}2-\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})+\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\bigg{\}}\\ &\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime}-\pi_{2}+\pi_{2}^{\prime})+2\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})-2\bigg{)}\cdot(\pi_{2}^{\prime}-\pi_{2})+\\ &\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime}).\end{split}$ (6)
In particular, we can find that if $\bar{\pi}\equiv\bar{\pi}^{\prime}$ then $(v(\bar{\pi})-v(\bar{\pi}^{\prime}))^{\rm T}(\bar{\pi}-\bar{\pi}^{\prime})\equiv 0$, thus $\bar{\pi}\equiv\bar{\pi}^{\prime}$ is one solution that $\pi_{1}$ and $\pi_{2}$ has reached equilibrium.

### v2

@@
 Abstract
-Chain-of-thought (CoT) prompting has emerged as a powerful technique for enhancing the reasoning capabilities of Large Language Models (LLMs) on complex problems.
-Among CoT-related studies, self-consistency (Multi-path inference with answer filtering through voting) involves generating multiple reasoning paths using the CoT framework and then selecting the most frequently produced outputs standing out as a concise yet competitive approach.
-While self-consistency has indeed led to the improvements in LLM inference, the use of multi-path inference also escalates deployment costs.
-Therefore, maintaining the performance benefits of self-consistency inherited from multi-path inference while reducing the inference costs holds significant value.
-In this research, we conceptualize language decoding as a preference consensus game, constructing a bi-player gaming system within each local path, and introduce Nash Chain-of-Thought (Nash CoT).
-Specifically, for a given question, we leverage LLM to autonomously select the contextually relevant template and generate outputs guided by this template, aiming to reach Nash Equilibrium alongside normal generation in each path.
-This approach allows us to achieve comparable or improved performance compared to self-consistency while using fewer inference paths on various inference tasks, including Arabic reasoning, Commonsense Question answering, and Symbolic inference.
+Chain of thought (CoT) is a reasoning framework that can enhance the performance of Large Language Models (LLMs) on complex inference tasks.
+In particular, among various studies related to CoT, multi-path inference stands out as a simple yet effective improvement.
+However, there is no optimal setting for the number of inference paths.
+Therefore, we have to increase the number of inference paths to obtain better results, which in turn increases the inference cost.
+To address this limitation, we can utilize question-related role templates to guide LLMs into relevant roles, thereby increasing the possibility of correct inferences for each path and further reducing dependence on the number of inference paths while improving reasoning accuracy.
+However, placing LLMs into specific roles may reduce their reasoning diversity and performance on a few tasks where role dependence is low.
+To alleviate the excessive immersion of the LLM into a specific role, we propose Nash CoT by constructing a competitive system on each path that balances the generation from role-specific LLMs’ and the general LLMs’ generation, thereby ensuring both effective role adoption and diversity in LLM generation further maintaining the performance of multi-path inference while reducing the requirement of the number of inference paths.
+We evaluate Nash CoT across various inference tasks, including Arabic Reasoning, Commonsense Question Answering, and Symbolic Inference, achieving results that are comparable to or better than those of multi-path CoT with the equal number of inference paths.
 Nash CoT: Multi-Path Inference with Preference Equilibrium
-Ziqi Zhang1thanks:    Equal contribution.,  Cunxiang Wang [^1] footnotemark: 1,  Xiong Xiao2,  Yue Zhang1thanks:    Corresponding authors.,  Donglin Wang122footnotemark: 2 1 School of Engineering, Westlake University 2 University of Cambridge {zhangziqi, wangcunxiang, zhangyue, wangdonglin}@westlake.edu.cn
+Ziqi Zhang1thanks:    Equal contribution.,  Cunxiang Wang [^1] footnotemark: 1,  Xiong Xiao2,  Yue Zhang1thanks:    Corresponding authors.,  Donglin Wang122footnotemark: 2 1 School of Engineering, Westlake University 2 University of Cambridge, UK stevezhangz@163.com {wangcunxiang, zhangyue, wangdonglin}@westlake.edu.cn
 1 Introduction
-Large Language Models (LLMs) have revolutionized Natural Language Processing (NLP) (Ouyang et al., 2022; etc., 2023; Jiang et al., 2023; Brown et al., 2020b; OpenAI, 2024).
-In particular, leveraging human-designed instructions as input, LLM demonstrates superior inference performance across various types of simple reasoning tasks (Radford et al., 2019; Brown et al., 2020a).
-But, its performance remains variable in complex reasoning tasks (Rae et al., 2022).
-To enhance LLM’s inference capabilities on complex issues, we can employ a step-by-step inference approach, known as Chain-of-Thought (CoT) prompting (Wei et al., 2023).
-For instance, starting with a template like "Let’s think step by step", LLM first generates rationales and then arrives at a final prediction.
-This approach significantly improves LLM’s inference performance across tasks like Arabic, Symbol Inference, and CommonsenseQA.
-Figure 1: Demonstrations of Self-Consistency and Nash Chain-of-Thought (Nash CoT).
-(a) Self-Consistency entails inferring multiple paths and subsequently voting the most frequently prediction.
-(b) In Nash CoT, inference occurs multiple times, with each path generating two responses, but only one reached Preference Equilibrium is sustained.
-Ultimately, the answer sampled is the one that reaches preference equilibrium.
-Subsequently, the impressive performance of CoT on complex inference tasks has spurred new developments in this direction (Wang et al., 2023; Wei et al., 2023; Jin and Lu, 2023; Zhang et al., 2022; Shi et al., 2022).
-Among these developments, self-consistency (Wang et al., 2023) emerges as a competitive CoT approach.
-It leverages uncertainty from multiple inference paths from a single LLM, and ranking generated answers by frequency can significantly enhance LLM’s inference performance.
-This approach significantly improves the performance of GPT-3 utilizing zero-shot CoT without any options for parameter tuning.
-Meanwhile, experimental results indicate that inference performance improves with the number of generated samples (Wang et al., 2023), suggesting that the potential of LLM’s inference capabilities has not yet to be fully exploited.
-Although self-consistency is straightforward to implement and requires no additional turning, it has a significant drawback: substantially higher inference costs compared to directly utilizing CoT.
-On the other hand, the performance improvements resulting from self-consistency imply that relying solely on single-path inference cannot fully harness the inference capabilities of LLM, and multi-path inference encompasses the potential correct answers.
-Therefore, it’s necessary to maintain this strategy.
-Meanwhile, numerous practical applications indicate that when a LLM is given the appropriate templates, it can perform specific tasks more proficiently.
-Hence, an intuitive strategy to reduce the number of inference paths in self-consistency is to use templates to guide the LLM to correctly infer each path.
+Large Language Models (LLMs) have profoundly revolutionized the field of Natural Language Processing (NLP) (Ouyang et al., 2022; etc., 2023; Jiang et al., 2023; Brown et al., 2020b; OpenAI, 2024).
+Specifically, leveraging human-designed instructions as input, LLMs demonstrate superior inference performance across various types of simple reasoning tasks (Radford et al., 2019; Brown et al., 2020a).
+However, in complex tasks, direct reasoning LLMs do not yield good results (Rae et al., 2022).
+To improve LLMs’ inference performance on complex inference tasks, it is popular to adopt a step-by-step reasoning framework known as Chain-of-Thought (CoT) prompting (Wei et al., 2023).
+For example, when appending templates that can guide LLMs to perform step-by-step reasoning, such as ‘‘Let’s think step by step", to the end of given question.
+We can lead the LLMs to output question-related rationale and further obtain the answer, enabling LLMs to achieve better performance than zero-shot reasoning.
+Figure 1: Demonstrations of Nash Chain-of-Thought (Nash CoT).
+As shown in this figure, Nash CoT can be divided into three main steps.
+Step 1 involves bringing the LLM into a template-related role.
+Step 2 utilizes the role-immersed LLM and LLM under normal conditions to collect model predictions separately.
+Step 3 filters the responses based on Nash Equilibrium.
+Subsequently, among various CoT-related improvements, modified by utilizing multi-path inference which is represented by Wang et al.
+(2023) is one of the most effective reasoning approaches.
+However, there is no theoretical proof identifying an optimal number of paths for multi-path reasoning.
+To ensure better inference performance, it is imperative to augment the inference paths.
+However, this might burden the inference budgets.
+In order to reduce the dependency of multi-path inference on the number of inference paths while further enhancing its effectiveness, we can cast LLMs into a relevant role for a specific question.
+For instance, we can utilize a template such as "You’re a mathematician, and $\cdots$" to cast the LLM in the role of a mathematician.
+This approach can improve the accuracy of each path’s answer generated by the LLM, thereby reducing the multi-path CoT’s reliance on the number of paths, potentially improving the LLM’s reasoning.
+However, excessively casting the LLM into a specific role using templates may compromise its robustness (diversity), thereby adversely affecting its performance in solving problems with minimal role dependency.
+To address this problem, we can introduce a certain degree of randomness into the reasoning process of the LLM.
+This can mitigate the constraints imposed by excessive role casting on the reasoning performance, ultimately maintaining the role-immersed multi-path CoT’s reasoning performance in these questions.
+Driven by this motivation, we construct a gaming system on each path, where the LLM in the role-playing context and the LLM in its normal state act as two distinct players.
+When this gaming system formed by these two players reaches a Nash Equilibrium, the reasoning under that path can balance the preferences of both role-playing with that of the LLM’s normal state.
+Furthermore, we can choose the response with the most Nash Equilibrium hits through voting to ensure the performance of LLM reasoning.
 Figure 2: General Performance Comparison. We compare the average performance of, zero-shot, and zero-shot CoT self-consistency (20 Paths) with our Nash CoT (10 Paths) on Mistral-Instruct and GLM4.
-To achieve this goal, we use the contextual information required by the question as a template to guide the inference in each path.
-This approach increases the probability that the LLM can correctly solve the question in each path, thereby potentially reducing the number of inference paths needed for multi-path inference.
-Meanwhile, to alleviate overconfidence in template-guided generation, we develop a bi-player gaming system that introduces the Nash Equilibrium of preference of strategy (defined as Preference Equilibrium) in multi-path inference.
-This system samples generations that balance the preferences of both template-guided and normal generation, ensuring the output robustly matches the context of the given question while sustaining some moderate generation pattern.
-Subsequently, we combine multi-path inference with Preference Equilibrium to propose Nash CoT.
-We present the comparison between Self-Consistency and Nash CoT in Figure 1.
-We conduct experiments with two local deployed LLMs-Mistral-Instruct (Jiang et al., 2023) and GLM4 (Zeng et al., 2022; Du et al., 2022) on various inference tasks, including Arabic Reasoning (Koncel-Kedziorski et al., 2015; Hosseini et al., 2014) and Symbolic Reasoning (Wei et al., 2023) and Commonsense Reasoning (Talmor et al., 2019; Geva et al., 2021).
-As shown in Figure 2, Nash CoT has achieved similar or even better performance against self-consistency with fewer inference paths.
-Meanwhile, as shown in Figure 3, Nash CoT has significantly reduced the inference cost by up to 50% on local deployed LLMs.
-To our knowledge, we are the first to introduce Preference Equilibrium into multi-path inference with the aim of harmonizing text generation guided by the optimal template with that generated by the model’s default state.
-It can reduce the number of paths needed to achieve good results in multi-path inference.
-And we are also the first to integrate Preference Equilibrium into multi-path inference and propose Nash CoT.
-It can achieve similar or even superior performance on various reasoning tasks while requiring only half costs compared to self-consistency.
-[^1] We will release our code at ¡https://github.com/stevezhangzA/nash-chain-of-thought¿.
+We conduct experiments with two local deployed LLMs that include Mistral-Instruct (Jiang et al., 2023) and GLM-4 (Zeng et al., 2022; Du et al., 2022) on various inference tasks, including Arabic Reasoning (Koncel-Kedziorski et al., 2015; Hosseini et al., 2014), Symbolic Reasoning (Wei et al., 2023) and Commonsense Reasoning (Talmor et al., 2019; Geva et al., 2021).
+As shown in Figure 2, Nash CoT can achieve similar or even better performance against self-consistency with fewer inference paths.
+Meanwhile, as shown in Figure 3, Nash CoT can significantly reduce the inference cost by up to 50% on locally deployed LLMs.
+To summarize, our contribution can be summarized as follows:
+  • To our knowledge, we are the first to introduce the concept of Nash Equilibrium into multi-path CoT, with the objective of balancing text generation guided by the role-immersed LLM with that produced by the LLM’s default state.
+This innovation enables us to minimize the number of paths required to achieve good results in multi-path inference, while simultaneously preserving the multi-path CoT’s inherent performance advantages.
+  • We evaluated the performance of Nash CoT on a wide range of reasoning tasks.
+Nash CoT can achieve similar results to self-consistency in Arabic Reasoning, Symbolic Reasoning, and Commonsense Question Answering tasks, given the same number of reasoning paths.
+[^1] We have released our code at ¡https://github.com/stevezhangzA/nash-chain-of-thought¿.
 2 Related Work
 Chain-of-Thought (CoT).
-There are three majority kinds of CoT approaches.
-Zero-shot CoT that Prompts LLM with simple yet instructions to guide LLM generate step by step (Kojima et al., 2023).
-Manual CoT that initialized by randomly sampled several cases from dataset or designed by human, and followed by utilizing these cases as demonstration to guided LLM generate follow the manner of demonstration (Wei et al., 2023), however, such methods can be biased if the demonstration can represent real distribution.
-Automatic (or batch) CoT (Zhang et al., 2022) first sample cases which have the most representative sentence embedding in each clusters, followed by inference with the same manner as manual CoT.
-Self-Consistency (Wang et al., 2023) showcases strong performance in vast benchmarks.
+There are three majority types of CoT approaches.
+The first is zero-shot CoT, which involves prompting the LLM with simple instructions to guide its generation step by step (Kojima et al., 2023).
+The second type is Manual CoT, which begins with sampling several cases from a dataset or manually designing templates.
+These cases are then used as demonstrations to guide the LLM in generating responses (Wei et al., 2023).
+However, such methods can be biased if the demonstrations aren’t sufficient to cover the optimal distribution.
+Automatic CoT (Zhang et al., 2022) first clusters the dataset, then samples the most representative sentence from each cluster as the demonstrations and finally uses these demonstrations to guide LLM inference.
+On the other hand, self-consistency (Wang et al., 2023) showcases strong performance in vast benchmarks.
 Apart from its impact on inference performance, self-consistency also boasts scalability as a key advantage.
-It seamlessly integrates with different approaches, such as tabular consistency of transformations (tab-CoT) Jin and Lu (2023), making it adaptable and versatile for various applications.
-Despite self-consistency has improved LLM’s performance on Arabic benchmarks.
-Correspondingly,self-consistency should have to inference multi-times, thus it burdens the deploy budgets.
-One way to address this limitation is by initially sampling multiple inference paths and then fine-tuning using the highest frequency path.
+It can seamlessly integrate with different approaches, such as tabular chain-of-thought (Tab-CoT) Jin and Lu (2023), making it adaptable and versatile for various applications.
+Furthermore, despite self-consistency can improve LLM’s performance on Arabic benchmarks, self-consistency must be inferred multiple times, burdening deployment budgets.
+One way to overcome this limitation is to infer multiple paths and fine-tune them based on the most frequent path.
 Specifically, Huang et al.
-(2022) suggest that gathering inferences from multiple paths and sampling the most frequent generation to fine-tune a smaller LLM can enhance the LLM’s inference performance.
+(2022) proposes that by gathering inferences from multiple paths and sampling the most frequent generation, we can enhance the inference performance of a smaller LLM.
 However, this approach still requires updating the LLM’s parameters, which is inefficient.
-Therefore, it is necessary to further explore inference methods to maintain the performance of self-consistency while reducing the number of inference paths.
+Therefore, it is necessary to further develop inference methods to maintain the performance of self-consistency while reducing the number of multi-path CoT inference paths.
 Preference Optimization.
-The training policy with Reinforcement Learning (RL) to reflect preference, termed Reinforcement Learning with Human Feedback (RLFH), was initially introduced by Akrour et al.
+Training policy with Reinforcement Learning (RL) to reflect preference, termed Reinforcement Learning with Human Feedback (RLFH), is initially introduced by Akrour et al.
 (2011) and has since undergone consistent improvement and refinement by Cheng et al.
@@
 (2012).
-This approach has been widely applied to adjust Large Language Models’ (LLMs) parameters to align with human preferences (Ouyang et al., 2022; Jiang et al., 2023; etc., 2023).
+This approach has been widely applied to adjust the parameters of LLMs in order to align them with human preferences (Ouyang et al., 2022; Jiang et al., 2023; etc., 2023).
 Recently, a new approach called Direct Optimizing from Preference (DPO) has been proposed by Rafailov et al.
-(2023), aiming to directly adjust LLMs to reflect human preferences without requiring a reward model and RL algorithm.
+(2023), aiming to directly adjust LLMs to reflect human preferences without requiring a reward model.
 Additionally, Munos et al.
-(2023) proposed combining DPO with Nash equilibrium to ensure convergence of the last iterated policy.
-Our study also utilizes the concept of equilibrium in preference model, but the main difference is that we utilize preference equilibrium as a standard to pick up the most preferred answer instead optimizing the LLM’s parameters.
+(2023) proposes combining DPO with Nash Equilibrium to ensure the convergence of the last iterated policy.
+Despite our study also utilizing the concept of Nash Equilibrium in the Preference model, the main difference is that we utilize Nash Equilibrium as a standard to pick up the answer instead of optimizing the LLM’s parameters.
 3 Preference Equilibrium in mini-batch inference
-Self-consistency has shown that the inference of LLMs under a single path may not represent their full capabilities.
-By simply conducting multiple inferences and filtering answers through voting, it is possible to achieve more accurate results.
-However, multi-path inference lacks a strong theoretical foundation to determine the optimal number of inference paths, potentially leading to much more computational resource consumption.
-To reduce the number of paths required for multi-path inference, we utilize the concept of Nash Equilibrium to locally construct a binary game system in multi-path inference.
-Specifically, the preference of each valid inference path of the LLM needs to achieve Nash equilibrium with the preferences of the generation guided by the template.
-This approach increases the probability of each path correctly answering the question while maintaining a certain level of robustness, thereby reducing the number of inference paths required by self-consistency.
+The experimental results of self-consistency indicate that multi-path reasoning can unleash the reasoning capability of LLMs.
+However, multi-path inference lacks a theoretical foundation for determining the optimal number of inference paths, requiring us to increase the number of paths to ensure performance, and further increase the computational resource consumption.
+To reduce the number of paths required for multi-path inference, we utilize the concept of Nash Equilibrium to locally construct a binary gaming system in multi-path inference.
+Specifically, the preference of each valid inference path of the LLM needs to achieve Nash equilibrium with the preferences of the generation guided by the role-immersed LLM.
+This approach increases the possibility of each path correctly answering the question while maintaining a certain level of robustness (diversity), thereby reducing the number of inference paths required by self-consistency.
 Preference Model.
-Given text input $x$, and the sampled predictions (answers) $y_{1}$, $y_{2}$, we first define $y_{1}$ prefers $y_{2}$ as Equation 1:
-$\begin{split}\mathcal{P}(y_{1}\prec y_{2}|x):={\rm sign}(r_{\theta}(y_{1}|x))-{\rm sign}(r_{\theta}(y_{2}|x)),\end{split}$ (1)
-[…]
-Player Templates: Templates for our preference model that are shown the structure: id(player): description.
-Mathematician: You are a mathematician, you excel at analyzing problems from a mathematical logical perspective and arrive at conclusions that align with your values.
+Given text input $x$, and the sampled predictions (answers) $y_{1}$, $y_{2}$, and Bradley-Terry reward model $r_{\theta}$ that has been deliberated by Chen et al.
+We first define $y_{1}$ is preferred over $y_{2}$ as Equation 1:
+$\begin{split}\mathcal{P}(y_{2}\prec y_{1}|x):={\sigma}(r_{\theta}(y_{1}|x)-r_{\theta}(y_{2}|x)),\end{split}$ (1)
+where $\mathcal{P}$ is preference model reflects the preference when inputting the pairs $(y_{1},y_{2})$.
+Player Templates : Role templates for our LLM that are shown the structure: {id(player): description.} Mathematician: You are a mathematician, you excel at analyzing problems from a mathematical logical perspective and arrive at conclusions that align with your values.
 Literary scholar: You are a literary scholar who has read a vast array of literary works.
@@
 $\cdots$ (other cases have been appended to the Appendix.)
-Additionally, drawing from the definition from Munos et al., we define one policy as more preferred over another as:
-$\begin{split}\mathcal{P}(\pi_{1}\prec\pi_{2}):=\mathbb{E}_{{}^{y_{1}\sim\pi_{1}(\cdot|x)}_{y_{2}\sim\pi_{2}(\cdot|x)}}[\mathcal{P}(y_{1}\prec y_{2}|x)]\end{split}$ (2)
+Furthermore, drawing from the definition in (Munos et al., 2023), we define one policy (LLM for decision making) $\pi_{1}$ as more preferred over another policy $\pi_{2}$ as:
+$\begin{split}\mathcal{P}(\pi_{2}\prec\pi_{1}):=\mathbb{E}_{{}^{y_{1}\sim\pi_{1}(\cdot|x)}_{y_{2}\sim\pi_{2}(\cdot|x)}}\bigg{[}\mathcal{P}(y_{2}\prec y_{1}|x)\bigg{]},\end{split}$ (2)
+subsequently, we imply the necessity and existence of Nash Equilibrium in the Preference model (shown in Equation 1).
 Preference Equilibrium.
-We aim to select the best template (we provide several cases in above example 3 for the current problem, thus facilitating the large model in problem-solving, correspondingly reduces the required num of inference paths.
-However, this may lead to some issues: 1) If the template is incorrectly chosen, it may cause the agent to generate answers outside the range of correct responses corresponding to the current problem, resulting in erroneous replies.
-2) The large model may excessively generate context-dependent responses, affecting its robustness.
-To address these issues, we build a local bi-player gaming system that the preference of template guided LLM (player 1) over normal status of LLM (player 2) is the pay-off of template guided LLM, vice visa.
-If player 1 and player 2 reaches Nash Equilibrium  [^2] Preference Equilibrium leverages the concepts of Nash Equilibrium.
+We first discuss why we should construct a gaming system: To increase the accuracy of each inference path, we can select the most suitable template (example 3) for the given question, immersing the LLM into a template-related role.
+This approach helps LLMs to solve problems more effectively, thereby reducing the number of inference paths needed.
+However, this may lead to some issues: If the LLM is excessively immersed in a specific role, it may reduce the diversity of LLM’s generation, and further reduce LLM’s accuracy in solving tasks that have low role dependence.
+To address these problems, we locally build a bi-player gaming system that the preference of role-immersed LLM (player 1) over the normal status of LLM (player 2) is the pay-off of role-immersed LLM, vice visa.
+If player 1 and player 2 reach Nash Equilibrium  [^2] Preference Equilibrium leverages the concepts of Nash Equilibrium.
 Nash equilibrium is proposed by John Nash.
-It is a concept solution where, assuming each participant knows the equilibrium strategies of the other participants, no participant can benefit by changing their own strategy., then the strategy can match the preference of both player 1 and player 2.
-Subsequently, we define the status that player 1 and player 2 reach Nash Equilibrium as Preference Equilibrium (Definition 1).
+It is a concept solution where, assuming each participant knows the equilibrium strategies of the other participants, no participant can benefit by changing their own strategy., then the generation can balance the preference of both player 1 and player 2.
+Definition 3.1 (Preference Equilibrium).
+Given any two policys (players) $\pi_{1}$ and $\pi_{2}$ along with the pay-off of $\pi_{1}$ and $\pi_{2}$ as $\mathcal{P}(\pi_{1}\prec\pi_{2})$ and $\mathcal{P}(\pi_{2}\prec\pi_{1})$, we define the Nash Equilibrium of $\pi_{1}$ and $\pi_{2}$ as Preference Equilibrium.
+Subsequently, we define the status that player 1 and player 2 reach Nash Equilibrium as Preference Equilibrium (Definition 3.1).
 Meanwhile, in Theorem 3.1, we prove the existence of Nash Equilibrium in this system.
-Specifically, the strategy of player 1 equal to player 2 is one solution that this system has reached Preference Equilibrium.
+In particular, the strategy of player 1 equal to player 2 is one solution that this system has reached Preference Equilibrium.
 Theorem 3.1 (Existence of Preference Equilibrium).
 Given any two policy (player) $\pi_{1}$ and $\pi_{2}$ within the gaming system defined in Definition 1, where $\pi\in\Pi$.
-$\pi_{1}\equiv\pi_{2}$denotes a solution where the gaming system reaches Nash Equilibrium.
-Proof of Theorem 3.1 see Appendix E.
+$\pi_{1}\equiv\pi_{2}$ is a solution where the gaming system reaches Nash Equilibrium.
+Proof of Theorem 3.1 see Appendix F.
 Meanings of the existence of Preference Equilibrium.
-Theorem 3.1 proves the existence of an Nash Equilibrium between the template guided LLM and the normal status of LLM.
-When reaching Preference Equilibrium, the preference of decisions made by the template guided LLM are aligned with those made by the LLM under normal status.
-Meanwhile, the preference of template guided LLM generation is much more closed to the requirement of quesitons’ context, while those of the normal status LLM are predominantly based on its parameters which is much more robust than template guided LLM.
-Therefore, this equilibrium can balance the requirement of contextual information and robustness of the model generation during problem-solving.
-Notabaly, $\pi_{1}\equiv\pi_{2}$ means their outputs are also likely to be equal.
-This insight forms a fundamental basis for piratically implementing Nash CoT.
+Theorem 3.1 proves the existence of a Nash Equilibrium between the role-immersed LLM and the normal status of LLM generation.
+Therefore, when reaching Preference Equilibrium, the preference of generations made by the role-immersed LLM is aligned with those made by the LLM without templates.
+Meanwhile, the preference for role-immersed LLM generation is much closer to the selected role-immersed context, while those of the normal status LLM are predominantly based on its parameters which are much more diverse than role-immersed LLM generation.
+Therefore, Preference Equilibrium can balance the requirement of role-immersed LLM generation and the diversity of the LLM generation in each path.
+In particular, $\pi_{1}\equiv\pi_{2}$ means their outputs are also likely to be equal.
+This insight forms a fundamental basis for practically implementing Nash CoT.
 3.1 Mini-batch inference with Preference Equilibrium
-Subsequently, based on the concept of Preference Equilibrium, we conceptualize a Mini-batch inference (shown in Figure 1) as a bi-player gaming system.
-This approach aims to achieve better inference compared to direct inference, while still preserving some of the inherent randomness of standard inference methods.
-To begin with proposing this system, we first define $x^{t}$ as the template of zero-shot CoT, $\{x^{c}_{0},x^{c}_{1},\cdots,x^{c}_{n}\}$ (we provided several cases in Player Templates) as the candidates template for template guided generation.
-Meanwhile, in this system, the template can to be chosen by a reward model $r_{\theta}$.
-[…]
-4 Nash Chain of Thought (Nash CoT)
+Subsequently, based on Preference Equilibrium, we conceptualize a mini-batch inference (shown in Step 2 of Figure 1) as a bi-player gaming system.
+This approach aims to maintain the performance benefits of multi-path inference while retaining some of the inherent randomness (diversity) of standard inference methods.
+Before introducing mini-batch inference, we first define $x^{t}$ as the template of zero-shot CoT.
+Meanwhile, we leverage LLMs to select a player template from the candidate set $\{x^{c}_{0},x^{c}_{1},\cdots,x^{c}_{n}\}$ for role-immersed generation, guided by Equation 1.
 Algorithm 1  Nash CoT (Answer Gathering)
@@
 12:  Return $\tau$
-Nash CoT can be seen as an extension of Mini-batch inference with Preference Equilibrium, implementing multiple Mini-batch inferences to enhance performance.
-This approach is influenced by experimental results from self-consistency, which suggest that increasing the number of paths can improve inference accuracy.
-Meanwhile, the reasoning process for each question is divided into two stages: Answer Gathering and Answer Filtering.
+[…]
+Based on the mini-batch inference, we further introduce Nash CoT in the next chapter.
+(In particular, in the current framework, successful inference cannot always be guaranteed.
+For instance, $y^{*}$ may not always in $[y_{1},y_{2}]$.
+We will address this problem in the following chapters.)
+4 Nash Chain of Thought (Nash CoT)
+Nash CoT can be regarded as an extension of mini-batch inference with Preference Equilibrium, implementing multiple mini-batch inferences to further enhance performance.
+This approach is inspired by self-consistency experiments, which suggest that increasing the number of paths enhances LLM’s inference accuracy.
+Specifically, Nash CoT’s reasoning process for each question can be divided into two stages: Answer Gathering and Answer Filtering.
 Answer Gathering.
-When generating candidate answers, the process predominantly involves two types of loops: Mini-batch Loops ($\rm n_{\rm mini}$): In Chapter 3.1, we discussed the implementation of mini-batch inference with Preference Equilibrium.
-As shown in Algorithm 1, this process involves searching for template-guided generations within two rounds of generation ($[y_{1},y_{2}]$).
-We refer to the times of these two predictions as the $\rm n_{\rm mini}$.
-Moreover, to mitigate the impact brought from low-frequency predictions, we introduce iterating $\rm n_{\rm mini}$ multiple times.
-This leads us to another type of loop: Outer Loops ($\rm n_{\rm outer}$): This loop resembles the concept of multi-path in self-consistency.
-After completing loop $\rm n_{\rm outer}$, we filter the generated answers and retain the answer that reaches equilibrium most frequently (shown in Algorithm 2), as the preferred answer.
+When generating candidate answers, the process predominantly involves two types of loops: mini-batch Loops ($\rm n_{\rm mini}$): As shown in Algorithm 1, this process involves searching for role-immersed generations within two rounds of generation $[y_{1},y_{2}]$.
+We denote the times of these two predictions as the $\rm n_{\rm mini}$.
+Moreover, to mitigate the impact brought by low-frequency predictions, we introduce iterating $\rm n_{\rm mini}$ multiple times.
+This leads us to another type of loop: Outer Loops ($\rm n_{\rm outer}$): This loop resembles the concept of multi-path inference that iterates $\rm n_{\rm mini}$ multiple times.
+After completing loop $\rm n_{\rm outer}$, we filter the generated answers and retain the answer that reaches Preference Equilibrium most frequently (shown in Algorithm 2), as the predicted answer.
 Core LLM Methods SingleEQ AddSub MultiArith GSM8K AQuA SVAMP Avg.
-Mistral-Instruct (7B) zero-shot 15.3 $\pm$ 0.8 12.0 $\pm$ 2.8 3.3 $\pm$ 1.3 2.7 $\pm$ 2.0 20.8 $\pm$ 1.5 7.7 $\pm$ 2.0 10.3$\pm$ 6.5
+Mistral-Instruct (7B) zero-shot 15.3$\pm$0.8 12.0$\pm$2.8 3.3$\pm$1.3 2.7$\pm$2.0 20.8$\pm$1.5 7.7$\pm$2.0 10.3$\pm$7
 zero-shot CoT 76.0$\pm$0.8 82.5$\pm$2.0 75.4$\pm$6.1 44.3$\pm$ 4.0 27.9$\pm$2.3 63.4$\pm$6.9 61.6$\pm$19
@@
 Nash CoT (10 Paths) 81.4$\pm$0.8 86.3$\pm$6.0 86.3$\pm$4.7 55.7$\pm$5.8 39.9$\pm$5.4 77.0$\pm$3.5 71.1$\pm$17
-GLM4-chat (9B) zero-shot 1.1 $\pm$ 1.5 1.1 $\pm$ 1.5 12.6 $\pm$ 3.9 12.0 $\pm$ 2.0 22.4 $\pm$ 4.1 4.4 $\pm$ 2.8 8.9$\pm$7.6
-zero-shot CoT 90.7 $\pm$ 1.5 90.7 $\pm$ 1.5 98.4 $\pm$ 1.3 80.9 $\pm$ 2.8 20.8 $\pm$ 3.1 86.9 $\pm$ 3.5 78.1$\pm$26.1
-self-consistency (20 Paths) 92.3 $\pm$ 2.0 90.2 $\pm$ 2.3 98.4 $\pm$ 2.3 89.3 $\pm$ 0.2 20.8 $\pm$ 3.1 91.5 $\pm$ 1.1 80.4$\pm$26.8
-Nash CoT (10 Paths) 91.3 $\pm$ 0.8 90.2 $\pm$ 2.7 96.7 $\pm$3.3 80.3$\pm$ 1.3 20.8$\pm$3.1 88.0 $\pm$2.0 77.9$\pm$26.1
-Table 1: Experimental results on arithmetic reasoning benchmarks.
+GLM4-chat (9B) zero-shot 1.1$\pm$1.5 1.1$\pm$1.5 12.6$\pm$3.9 12.0$\pm$2.0 22.4$\pm$4.1 4.4$\pm$2.8 8.9$\pm$8
+zero-shot CoT 90.7$\pm$1.5 90.7$\pm$1.5 98.4$\pm$1.3 80.9$\pm$2.8 20.8$\pm$ 3.1 86.9$\pm$3.5 78.1$\pm$26
+self-consistency (20 Paths) 92.3$\pm$2.0 90.2$\pm$2.3 98.4$\pm$2.3 89.3$\pm$0.2 20.8$\pm$3.1 91.5$\pm$1.1 80.4$\pm$27
+Nash CoT (10 Paths) 91.3$\pm$0.8 90.2$\pm$2.7 96.7$\pm$3.3 80.3$\pm$1.3 20.8$\pm$3.1 88.0$\pm$2.0 77.9$\pm$26
+Table 1: Experimental results on Arabic Reasoning benchmarks.
 We test Zero-Shot CoT and Nash CoT with the core LLM includes Mistral-Instruct (7B) and GLM4-chat (9B) on mathematical benchmarks including AddSub, MultiArith, SingleEQ, SVAMP, GSM8K, and AQuA.
 Nash CoT performs the best.
-Answer Filtering.
-In terms of answer filtering, as shown in Algorithm 2 we first count the most frequent prediction satisfy Preference equilibrium.
-Specifically, we count all $y^{*}$ satisfy $y^{*}\in[y_{1},y_{2}]$ and compute their frequency.
-Subsequently, we return the most frequent case.
-Otherwise, if is no cases satisfy $y^{*}\in[y_{1},y_{2}]$, we adopt the strategy of self-consistency by selecting the most frequent prediction among all generated answers.
 Algorithm 2  Nash CoT (Answer Filtering)
@@
 12:  end if
-Subsquently, we propose Nash CoT, which iterates through Algorithm 1 and Algorithm 2 to perform inference on all sampled questions, where $\tau$ represents the candidate answers from the Answer Gathering stage.
+Answer Filtering.
+In terms of answer filtering, as shown in Algorithm 2 we first count the most frequent prediction to satisfy Preference equilibrium.
+Specifically, we count all $y^{*}$ satisfy $y^{*}\in[y_{1},y_{2}]$ and compute their frequency.
+Subsequently, we return the most frequent answer.
+Otherwise, if no cases satisfy $y^{*}\in[y_{1},y_{2}]$, we adapt answer filtering by selecting the most frequent prediction among all generated answers.
 […]
-Practical Implementation of reward model $r_{\theta}$.
-In the process of practical implementation, we do not explicitly train a reward model $r_{\theta}$ to confine the player template $x^{c}$ (we have provided cases in Player Templates) using Equation 1.
-Instead, we directly use the preference template (shown in Preference Template) to guide the LLM in determining the most suitable player template for a given question.
-For example, when presented with a coin flip question as shown in Figure 4, we fill the Preference Template with given question and player templates.
+Practical Implementation of Preference model $\mathcal{P}$.
+In the process of practical implementation, we do not explicitly train a reward model $r_{\theta}$ to confine the player template $x^{c}$ utilizing Equation 1.
+Instead, we directly utilize the template (shown in Preference Template) to guide the LLM in determining the most suitable player template for a given question.
+For example, when presented with a coin flip question as shown in Figure 4, we fill the Preference Template with the given question and Player Templates.
 This filled template is then input into the LLM to provide the id of the most suitable player template from the available options.
-In particular, we believe it’s effective, this is because most of baselines we selected have been turned to reflect human preference, thus we believe the selected LLM can be directly utilized as the preference model to point out the most preferred option among candidate options.
+In particular, we consider utilizing LLM to search templates to be effective, this is because most of the baselines we selected have been turned to reflect human preference, thus LLM can be directly utilized as the preference model to point out the most preferred option among candidate options (we provide cases in section A).
+Subsequently, we propose Nash CoT, which iterates through Algorithm 1 and Algorithm 2 to perform inference on all sampled questions, where $\tau$ represents the candidate answers from the Answer Gathering stage.
 5 Experiments
@@
 self-consistency (20 Paths) 21.9$\pm$4.7 0.0$\pm$0.0 38.8$\pm$0.8 47.0$\pm$1.5 26.9$\pm$18
-Nash CoT (10 Paths) 29.0$\pm$ 5.4 0.5 $\pm$ 0.8 44.8 $\pm$ 2.0 41.1 $\pm$ 1.2 28.9$\pm$ 17
-GLM4-chat (9B) zero-shot 27.3 $\pm$ 6.3 0.0 $\pm$ 0.0 38.8 $\pm$ 0.8 16.4 $\pm$ 2.3 20.7$\pm$14.3
-zero-shot CoT 87.4 $\pm$ 0.8 0.0 $\pm$ 0.0 37.7 $\pm$ 2.3 16.4 $\pm$ 4.8 35.4$\pm$32.9
-self-consistency (20 Paths) 98.9$\pm$1.5 0.0$\pm$0.0 37.7$\pm$2.3 16.4$\pm$4.8 38.3$\pm$37.5
-Nash CoT (10 Paths) 93.4$\pm$2.7 0.0 $\pm$ 0.0 37.7$\pm$2.3 16.4$\pm$4.8 36.9$\pm$35.3
+Nash CoT (10 Paths) 29.0$\pm$5.4 0.5$\pm$0.08 44.8$\pm$2.0 41.1$\pm$1.2 28.9$\pm$17
+GLM4-chat (9B) zero-shot 27.3$\pm$6.3 0.0$\pm$0.0 38.8$\pm$0.8 16.4$\pm$2.3 20.7$\pm$14
+zero-shot CoT 87.4$\pm$0.8 0.0$\pm$0.0 37.7$\pm$2.3 16.4$\pm$4.8 35.4$\pm$33
+self-consistency (20 Paths) 98.9$\pm$1.5 0.0$\pm$0.0 37.7$\pm$2.3 16.4$\pm$4.8 38.3$\pm$38
+Nash CoT (10 Paths) 93.4$\pm$2.7 0.0$\pm$0.0 37.7$\pm$2.3 16.4$\pm$4.8 36.9$\pm$35
 Table 2: Experimental results on symbolic inference benchmarks.
-We test Zero-Shot CoT and Nash CoT with Mistral-Instruct (7B) and GLM4-chat (9B) on Symbolic QA benchmarks includes Coin-Flipping, Last Letters and Object Tracking.
+We test Zero-Shot CoT and Nash CoT with Mistral-Instruct (7B) and GLM4-chat (9B) on Symbolic QA benchmarks including Coin-Flipping, Last Letters, and Object Tracking.
 Among these baselines, Nash CoT performs the best.
@@
 Nash CoT (10 Paths) 56.8$\pm$2.0 69.4$\pm$4.7 63.1$\pm$6
-GLM4-chat (9B) zero-shot 56.8 $\pm$ 4.7 17.5$\pm$ 2.0 37.2$\pm$19.7
-zero-shot CoT 63.9$\pm$2.3 18.0$\pm$2.3 41.0$\pm$ 22.9
-self-consistency (20 Paths) 69.9 $\pm$3.3 18.0$\pm$2.3 44.0$\pm$26.0
-Nash CoT (10 Paths) 66.7 $\pm$ 0.8 18.0 $\pm$ 2.3 42.4$\pm$24.4
+GLM4-chat (9B) zero-shot 56.8$\pm$4.7 17.5$\pm$2.0 37.2$\pm$20
+zero-shot CoT 63.9$\pm$2.3 18.0$\pm$2.3 41.0$\pm$23
+self-consistency (20 Paths) 69.9$\pm$3.3 18.0$\pm$2.3 44.0$\pm$26
+Nash CoT (10 Paths) 66.7$\pm$0.8 18.0$\pm$2.3 42.4$\pm$24
 Table 3: Experimental results on Commonsense Reasoning.
 We test Zero-Shot CoT and Nash CoT with Mistral-Instruct (7B) and GLM4-chat (9B) on Commonsense Reasoning datasets includes StrategyQA and CommonsenseQA .
-The goal of our experiment is to 1) demonstrate the performance advantage and effectiveness of Nash CoT.
-2) shocase whether Nash CoT help reduce the overall inference time.
-In the following sections, we first introduce our experimental setup and then present the experimental results and analysis.
+The goal of our experiment is to 1) showcase the performance advantage and effectiveness of Nash CoT. 2) showcase whether Nash CoT helps reduce inference paths and inference time.
 Datasets.
 Our majority benchmarks are composed of three different kinds of inference tasks.
-1) arithmetic reasoning: SingleEq (Koncel-Kedziorski et al., 2015), AddSub (Hosseini et al., 2014), MultiArith (Roy and Roth, 2016), GSM8K (Cobbe et al., 2021), AQUA (Ling et al., 2017), and SVAMP (Patel et al., 2021).
+1) Arabic Reasoning: SingleEq (Koncel-Kedziorski et al., 2015), AddSub (Hosseini et al., 2014), MultiArith (Roy and Roth, 2016), GSM8K (Cobbe et al., 2021), AQUA (Ling et al., 2017), and SVAMP (Patel et al., 2021).
 2) Symbolic Reasoning: Last Letters, Coin Flip (Wei et al., 2023), and Object Tracking, Bigbench Date.
@@
 LLMs.
-To validate that Nash CoT is a general CoT method, we selected different large models as test models, including Mistral-7B (Instruct) (Jiang et al., 2023), GLM4-9B-Chat Zeng et al.
+To validate that Nash CoT is a general CoT method, we selected different large models as test models, including Mistral-Instruct (7B) (Jiang et al., 2023), Chat-GLM4 (9B) Zeng et al.
 (2022); Du et al.
 (2022).
-In particular, all of these selected LLMs are turned via RL with human feedback (RLHF), and the difference between LLM turned with RLHF and the original foundation models have been detailed by Ouyang et al.
-(2022).
+In particular, all of these selected LLMs are turned via RLHF, and the difference between LLM turned with RLHF and the original foundation models have been detailed by Ouyang et al..
 Baselines.
-The preliminary baselines we utilized include zero-shot, zero-shot CoT (Wei et al., 2023), andself-consistency (Wang et al., 2023). We test these approach with freezed LLMs.
+The preliminary baselines we utilized include zero-shot, zero-shot CoT (Wei et al., 2023), and self-consistency (Wang et al., 2023). We test these approaches with frozen LLMs.
 Settings.
-Our evaluation on all selected tasks utilizes the same experimental settings bellow:
+Our evaluation of all selected tasks utilizes the same experimental settings below:
   • zero-shot and zero-shot CoT. We follow the method proposed by Wei et al. (2023) and use the original template (e.g., "Let’s think step by step") for evaluation.
@@
 We follow Wang et al.
-(2023) to evaluate the performance ofself-consistency with selected LLMs, utilizing the zero-shot CoT template.
+to evaluate the performance of self-consistency with selected LLMs, utilizing the zero-shot CoT template.
 Additionally, we set the number of inference paths to 20.
@@
 We set up $n_{\rm outer}$ as 3 and $n_{\rm mini}$ as 2, resulting in a total of $n_{\rm outer}\times(n_{\rm mini}+1)+1=10$ paths.
-Additionally, we have provided the player templates $x^{t}$ in Table 1 and the Appendix, meanwhile we utilizing the same CoT template $x^{c}$ as in zero-shot CoT.
-Additionally, all evaluations are conducted on the inference of 60 random sampled questions multi times. And we have provided the mean and standard error in all tables.
-Figure 3: We used GLM4-chat (9B) on the same type of GPU (A-100) to evaluate Nash CoT and self-consistency across selected tasks.
+Additionally, we have provided the Player Templates $x^{t}$ in Table 1 and the Appendix, meanwhile, we utilizing the same CoT template $x^{c}$ as in zero-shot CoT.
+Specifically, in Table 12 and 3, we set the total number of reasoning paths for Nash CoT to 10 and for self-consistency to 20.
+The primary goal is to demonstrate that Nash CoT can achieve the same or even better results with only half the number of reasoning paths compared to self-consistency.
+Additionally, all evaluations are conducted on the inference of 60 random sampled questions per seed. We have provided the mean and standard error in the majority tables.
+Figure 3: We used GLM4-chat (9B) on the same type of GPU (A100) to evaluate Nash CoT and self-consistency across selected tasks (60 questions per task).
 Nash CoT, employing a total of 10 paths, requires nearly half the time of self-consistency, which has 20 paths in total.
-5.1 Experimental Results
-Evaluated Scores.
-The majority experimental results are demonstrated in table 12 and 3.
-Nash CoT can improve Mistral-Instruct (7B) on almost all selected inference tasks, while showcasing similar performance to self-consistency with twice inference paths on GLM4-chat (9B).
-In particular, we have provided the total paths of Nash CoT that it only require the half of self-consistency, thus our claim in section 3 can be validated.
 Figure 4: We use Mistral-Instruct (7B) to examine the impact of loop numbers on the inference performance of the large language model.
@@
 We marked self-consistency with 20 paths using dashed lines, and some results of Nash CoT, with total paths close to 20, were marked with stars.
-When focusing on Mistral-Instruct (7B), Nash CoT has better performance on arithmetic and symbol inference tasks, showcasing its superiority performance on logic/math inference tasks.
+5.1 Experimental Results
+Models Settings Cnt and Prob Precalculus Inter-Alg Number Theory Avg.
+Mistral-Instruct (7B) zero-shot-cot 40.9 16.4 16.4 24.6 24.6
+Nash CoT 64.8 38.5 36.9 27.9 42.0
+self-consistency 72.1 34.4 34.4 21.3 40.6
+GLM4 (large) zero-shot-cot 83.6 90.2 85.2 91.9 87.7
+Nash CoT 95.1 90.2 98.3 98.3 95.5
+self-consistency 93.3 90.2 95.1 100.0 94.7
+Table 4: Equal Path Evaluation.
+This experiment is conducted to verify the performance of Nash CoT when the number of inference paths is set equal to self-consistency.
+In particular, Cnt and Prob denote Counting and Probability, while Inter-Alg denotes Intermediate Algebra.
+Evaluated Scores.
+The majority experimental results are demonstrated in Table 12 and 3.
+Nash CoT can improve Mistral-Instruct (7B) on almost all selected inference tasks while showcasing similar performance to self-consistency with twice inference paths on GLM4-chat (9B).
+In particular, we have provided the total paths of Nash CoT that it only require half of self-consistency, thus our claim in section 3 can be validated.
+When focusing on Mistral-Instruct (7B), Nash CoT has better performance on Arabic and Symbolic inference tasks, showcasing its superior performance on logic/math inference tasks.
 However, Nash CoT does not showcase improved performance in Commonsense Question Answering tasks.
 We argue that this is because Commonsense Question Answering tasks are more diverse, and the player template can’t cover all topics.
-Therefore, the player template limits Nash CoT on commonsense question answering tasks.
-Importantly, we limit Nash CoT’s performance by utilizing only total 10 paths for inference in this section.
-However, additional experimental results in the ablation section show that Nash CoT outperforms self-consistency via increasing the inference loops, thus Nash CoT can outperform self-consistency.
+Therefore, the predefined player template limits Nash CoT on Commonsense Question Answering tasks.
+Importantly, we evaluate Nash CoT by utilizing only a total of 10 paths for inference in this section.
+However, additional experimental results in the ablation section show that Nash CoT outperforms self-consistency by increasing the inference loops.
+Meanwhile, we also evaluate by setting equal paths on much more challenging MATH Hendrycks et al.
+(2021), showcasing that Nash CoT can outperform self-consistency.
+Comparing Nash CoT with self-consistency with equal paths.
+To further evaluate the performance of Nash CoT, we utilize Mistral-Instruct (7B) while selecting one of state-of-the-art (SOTA) LLMs named GLM4 (large) (GLM et al., 2024), for comparison using an equal number of reasoning paths (10 paths).
+As shown in Table 4, Nash CoT outperforms self-consistency, which demonstrates the superiority of Nash CoT compared to self-consistency when the same number of path settings are set.
 Inference Time.
@@
 6 Ablation Study
-In order to further validate the effectiveness of Nash CoT, we conducted extensive ablations to answer the following questions: 1) What will happen when the number of inference paths for Nash CoT is further increased?
-Will Nash CoT eventually surpass self-consistency, and what is the relationship between the number of loops and performance?
-2) Does the template really improve the accuracy of path predictions, and what impact does it have on experimental performance?
+We further conduct extensive ablations to answer the following questions: 1) Is there a correlation between the performance of Nash CoT and the setting of the number of inference paths and can the performance of Nash CoT be improved to surpass self-consistency by adjusting the number of inference paths?
+2) Does the template improve the accuracy of path predictions, and what impact does it have on experimental performance?
+3) Is the Nash Equilibrium the reason for the performance improvement observed in Nash CoT?
 As the number of inference paths increases, Nash CoT can obviously surpass self-consistency with fewer inference paths.
-To address question 1), we selected Mistral-Instruct (7B) and conducted evaluation on three different reasoning tasks, adjusting the $N_{\rm mini}$ and $N_{\rm outer}$.
-As shown in Figure 4, as the number of loops increases, Nash CoT has a high probability of significantly outperforming self-consistency with fewer paths.
-However, different from self-consistency, the experimental results of Nash CoT do not show a monotonic (linear) relationship with the total number of total paths.
+To address question 1), we selected Mistral-Instruct (7B) and evaluated three different reasoning tasks, adjusting the $N_{\rm mini}$ and $N_{\rm outer}$.
+As shown in Figure 4, as the number of loops increases, Nash CoT has a high possibility of outperforming self-consistency with fewer paths.
+However, different from self-consistency, the experimental results of Nash CoT do not show a monotonic (linear) relationship with the total number of paths.
 This indicates that there is a significant difference between Nash CoT and self-consistency.
@@
 Results are shown in Table 5, showing an approximately 9.2% decrement in GSM8K and 6.2% decrement in SVAMP.
-Therefore, the performance of Nash CoT is impacted by the Player Template.
+Therefore, the performance of Nash CoT is impacted by the Player Templates.
 GSM8K AQuA SVAMP
@@
 Table 5: Performance decreasing. We remove the mathematics from Player Templates and test Nash CoT on selected Arabic Reasoning tasks.
+As shown in Table 6, we used three role templates with varying degrees of relevance to mathematical problems: mathematician, student, and poker player, to test the performance of Mistral-Instruct (7B) on GSM8K.
+Role templates with higher relevance to mathematical problems can bring more improvements to Mistral, further confirming the rationality of enhancing reasoning performance by substituting LLM into roles using templates.
+mathematician student poker player
+57.37 52.45 47.54
+Table 6: Results on GSM8K. The templates for these roles are shown in Table 9 on the Appendix.
+Effectiveness of Nash Equilibrium.
+While Nash CoT has demonstrated effectiveness across various LLMs and benchmarks, however, the extent to which Nash CoT’s performance is driven by Nash Equilibrium remains to be determined.
+To further investigate, we compare the self-consistency integrated template (denoted as self-con.+temp (10 paths)) with Nash CoT (10 paths) using Mistral-Instruct (7B).
+As shown in Table 7, Nash CoT consistently outperforms self-con.+temp when applying the same mathematical templates across selected tasks.
+Settings AQuA SVAMP GSM8K AddSub Avg.
+self-con.+temp 31.7 71.6 53.6 85.8 60.6
+Nash CoT 39.9 77.0 55.7 86.3 64.7
+Table 7: Ablation of Nash Equilibrium. This experiment is conducted to verify the necessity of Nash Equilibrium in the performance of Nash CoT.
 7 Conclusion
-In this study, we proved the existence of Nash equilibrium in preference model, subsequently, we proposed a new CoT approach Nash CoT, and validated its performance on various inference benchmarks.
-Experimental results show that Nash CoT can perform equally or even better than self-consistency while only require half inference costs.
+We prove the existence of a Nash equilibrium in preferences between role-immersed and normal-status LLM generations, and develop a gaming system within each path of the multi-path CoT to introduce Nash CoT.
+Experimental results show that Nash CoT can perform equally or even better than self-consistency while requiring fewer paths.
 Limitations and Future Work
-Despite Nash CoT showcase competitive performance with only half of inference paths, it requires pre-defined template, thus it’s in-convenient to utilize Nash CoT in new emerging scenario, in the future we will develop a automatic approach to balance task feedback and template design.
+Since Nash CoT relies on predefined templates, its performance will decline in new scenarios due to limited adaptability in template selection.
+Future efforts will concentrate on automating the balancing of task feedback and template selection.
+Acknowledgement
+This work was supported by the National Science and Technology Innovation 2030 - Major Project (Grant No.2022ZD0208800), and NSFC General Program (Grant No.62176215).
+Meanwhile, we thank the reviewers for their valuable suggestions.
 Ethics Claims
-Despite LLM has showcased superiority performance on vast benchmarks, but pre-train or fine-tune a LLM requires numerous computing resources.
-Therefore, it’s crucial to study how to inference a LLM to reach the ceiling of its capacity.
-CoT is a ideal approach which has been proved that can obviously evaluate the performance of LLMs’ inference.
-Among that,self-consistency is one of the best CoT approach.
-Our method effectively reduce the inference times of multi-path inference, thereby reducing the deploy budgets of self-consistency.
+Pre-training or fine-tuning an LLM requires much more computing resources.
+Multi-path CoT is an ideal approach that has been proven to enhance the performance of LLMs’ inference.
+Meanwhile, Nash CoT effectively reduces the inference paths needed and times of multi-path inference.
 We believe our approach can further elevate the effectiveness of multi-path inference, thereby further improving the effectiveness of LLM.
-acl
 Appendix A Dataset
 Our majority dataset are composed of three different kinds of inference tasks.
-1) arithmetic reasoning: SingleEq (Koncel-Kedziorski et al., 2015), AddSub (Hosseini et al., 2014), MultiArith (Roy and Roth, 2016), GSM8K (Cobbe et al., 2021), AQUA (Ling et al., 2017), and SVAMP (Patel et al., 2021).
-2) symbolic reasoning: Last Letters, Coin Flip (Wei et al., 2023), and Object Tracking, Bigbench Date.
-3) commonsense question answering: CommonsenseQA Talmor et al.
-(2019) and StrategyQA Geva et al.
-(2021).
-For more details about the dataset please refer to Wang et al.
-(2023).
-Appendix B Uesage of LLM.
+  • Arabic Reasoning: SingleEq (Koncel-Kedziorski et al., 2015), AddSub (Hosseini et al., 2014), MultiArith (Roy and Roth, 2016), GSM8K (Cobbe et al., 2021), AQUA (Ling et al., 2017), SVAMP (Patel et al., 2021), and MATH dataset (Hendrycks et al., 2021).
+In particular, MATH dataset are process by our own, : When constructing the dataset, we identify the answers to each question in the MATH dataset through ground truth (GT) reasoning, retaining a certain number of questions for each category.
+When testing the LLM, we input the questions, collect the responses from the deployed LLM, and determine whether the LLM’s reasoning is correct by comparing whether the inferred results match the GT, examples are shown in Table 11.
+Meanwhile, the dataset or its url will be released at our codebase page.
+    Dataset Capacity
+    Algebra 778
+    Counting and Probability 291
+    Geometry 235
+    Intermediate Algebra 443
+    Number Theory 474
+    Prealgebra 618
+    Precalculus 133
+    Table 8: Data type and data capacity
+  • Symbolic Reasoning: Last Letters, Coin Flip (Wei et al., 2023), and Object Tracking, Bigbench Date.
+  • Commonsense Question Answering: CommonsenseQA Talmor et al. (2019) and StrategyQA Geva et al. (2021). For more details about the dataset please refer to Wang et al. (2023).
+Appendix B Experimental Settings
+In this section, we compare Nash CoT with self-consistency, focusing specifically on self-consistency as a representative method.
+As illustrated in Figure 1, Nash CoT operates with two distinct loops, where the final answer is derived from voting on the response that achieves Preference Equilibrium most frequently.
+In contrast, self-consistency determines the most frequent answer across all predictions.
+Figure 5: Comparison of Nash CoT and self-consistency. (right) Nash Chain of thought (Nash CoT). (left) self-consistency.
+Appendix C Usage of LLM.
 We utilize LLM to rectify grammar errors.
-Appendix C Computing Resources
+Appendix D Computing Resources
 Our experiments were run on a computer cluster with 32GB RAM, 4-Core CPU, and NVIDIA-A100 (80G, 32G)/NVIDIA-V100 (32G) GPU, Linux platform.
-Appendix D Source Code.
+Appendix E Source Code.
 We have provided source code for reference.
 Additionally, our code are based on https://github.com/amazon-science/auto-cot and refer to the coding manner from https://github.com/eureka-research/Eureka.
-Appendix E Proof of theorem 3.1.
+Appendix F Proof of theorem 3.1.
 Subsequently, we prove the existence of Nash equilibrium in this system.
@@
 $\begin{split}(v(\bar{\pi})-v(\bar{\pi}^{\prime}))^{\rm T}(\bar{\pi}-\bar{\pi}^{\prime})\leq 0,\end{split}$ (5)
-where $\bar{\pi}$ and $\bar{\pi}^{\prime}$ are any two given policy set. Subsequently, we can further darrive at the following relationships:
-$\begin{split}(v(\bar{\pi})-v(\bar{\pi}^{\prime}))^{\rm T}(\bar{\pi}-\bar{\pi}^{\prime})&=\begin{pmatrix}R(\pi_{1};\pi_{2})-R(\pi_{1}^{\prime};\pi_{2}^{\prime})\\ R(\pi_{2};\pi_{1})-R(\pi_{2}^{\prime};\pi_{1}^{\prime})\\ \end{pmatrix}\cdot\begin{pmatrix}\pi_{1}-\pi_{1}^{\prime},\pi_{2}-\pi_{2}^{\prime}\\ \end{pmatrix}\\ &=\bigg{(}R(\pi_{1};\pi_{2})-R(\pi_{1}^{\prime};\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime})+\bigg{(}R(\pi_{2};\pi_{1})-R(\pi_{2}^{\prime};\pi_{1}^{\prime})\bigg{)}\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime})+\bigg{\{}2-\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})+\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\bigg{\}}\\ &\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime}-\pi_{2}+\pi_{2}^{\prime})+2\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})-2\bigg{)}\cdot(\pi_{2}^{\prime}-\pi_{2})+\\ &\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime}).\end{split}$ (6)
+where $\bar{\pi}$ and $\bar{\pi}^{\prime}$ are any two given policy set. Subsequently, we can further arrive at the following relationships:
+$\begin{split}(v(\bar{\pi})-v(\bar{\pi}^{\prime}))^{\rm T}(\bar{\pi}-\bar{\pi}^{\prime})&=\begin{pmatrix}R(\pi_{1};\pi_{2})-R(\pi_{1}^{\prime};\pi_{2}^{\prime})\\ R(\pi_{2};\pi_{1})-R(\pi_{2}^{\prime};\pi_{1}^{\prime})\\ \end{pmatrix}\cdot\begin{pmatrix}\pi_{1}-\pi_{1}^{\prime},\pi_{2}-\pi_{2}^{\prime}\\ \end{pmatrix}\\ &=\bigg{(}R(\pi_{1};\pi_{2})-R(\pi_{1}^{\prime};\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime})+\bigg{(}R(\pi_{2};\pi_{1})-R(\pi_{2}^{\prime};\pi_{1}^{\prime})\bigg{)}\\ &\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime})+\bigg{\{}2-\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})+\\ &\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\bigg{\}}\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime}-\pi_{2}+\pi_{2}^{\prime})+2\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})-2\bigg{)}\cdot(\pi_{2}^{\prime}-\pi_{2})+\\ &\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime}).\end{split}$ (6)
 In particular, we can find that if $\bar{\pi}\equiv\bar{\pi}^{\prime}$ then $(v(\bar{\pi})-v(\bar{\pi}^{\prime}))^{\rm T}(\bar{\pi}-\bar{\pi}^{\prime})\equiv 0$, thus $\bar{\pi}\equiv\bar{\pi}^{\prime}$ is one solution that $\pi_{1}$ and $\pi_{2}$ has reached equilibrium.
+Appendix G Further clarify our settings
+Why in Table 12 and 3 do we set the number of pathways for Nash CoT to be half of self-consistency?
+If Nash CoT can surpass self-consistency with half the number of reasoning pathways, then it is easier to demonstrate the validity and efficiency of Nash CoT.
+Because that according to the experimental conclusion in Wang et al.
+that ‘‘The experimental effect of self-consistency increases as the number of pathways increases.", if Nash-CoT outperforms self-consistency with fewer paths, it indicates that Nash-CoT is better than self-consistency across all settings with an equal or greater number of paths.
+Which one performs better between Nash CoT and self-consistency when they have the same number of pathways?
+In our ablation experiments to test the impact of the number of pathways on Nash CoT, we found that by adjusting the number of pathways, Nash CoT can surpass self-consistency with fewer pathways on Aqua, CommensenseQA, and Object Tracking.
+Additionally, in Table 4, we used GLM4 (large), Mistral-Instruct (7B) to outperform self-consistency on the MATH dataset with the same number of pathways.
+Appendix H Case Study
+H.1 LLM as preference model
+In this experiment, we provided examples of LLM inferences to demonstrate that LLMs can select the most appropriate template based on the question.
+Specifically, in Table 9, we set up a series of templates for different roles, arranging them in different sequences and labeling them with numbers.
+These were input into the model, allowing it to rank the templates according to its preference for the question.
+We observe in Table 10 that the order given by the LLM corresponded to the performance strengths of different templates on the respective tasks, which proves that LLMs can directly serve as preference models to select the optimal template for a specific question.
+Player Description Correlation
+Mathematician You are a mathematician, You excel at analyzing problems from a Mathematical logical perspective and arrive at conclusions that align with your values. High
+Student You are a student, please answer the following questions. Medium
+Poker Player You are playing a poker game, please do your best to play poker games. Low
+Table 9: Player templates we utilized.
+Input Ranking
+We have four templates: 1.
+Mathematician: You are a mathematician, You excel at analyzing problems from a Mathematical logical perspective and arrive at conclusions that align with your values.
+2.
+Student: You are a student, please answer the following questions.
+3.
+Poker Player: You are playing a poker game, please do your best to play poker games.
+Please rank their correlations to a math problem, and directly output the sequential ranked numerical tag.
+1, 2, 3
+We have four templates: 1.
+Student: You are a student, please answer the following questions.
+2.
+Mathematician: You are a mathematician, You excel at analyzing problems from a Mathematical logical perspective and arrive at conclusions that align with your values.
+3.
+Poker Player: You are playing poker game, please do your best to play poker games.
+Please rank their correlations to math problem, and directly output the sequential ranked numerical tag.
+2, 1, 3
+Table 10: We provide cases for the ranking provided by LLM.
+The roles of Mathematician, Student, and Poker Player achieved scores of 57.37, 52.45, and 47.54, respectively, on GSM8K.
+The performance of these different roles is consistent with the model’s ranking.
+Meanwhile, we utilize ChatGPT to rank these templates based on their relevance to the topic of questions.
+Our inputs are shown in Table 10.
+The model’s ranking of different roles is consistent with their performance levels on the mathematical dataset.
+Meanwhile, changing the input order of the templates simultaneously does not affect the model’s ranking of the templates.
+Appendix I Format of our processed MATH dataset
+We provide several cases for readers to understand the format of our processed MATH dataset.
+Dataset Question Answer Capacity
+Algebra How many vertical asymptotes does the graph of $y=\frac{2}{x^{2}+x-6}$ have? 2 778
+Counting and Probability $n$ fair 6-sided dice are simultaneously rolled. The probability that exactly two of them show a number other than 1 is $\frac{25}{216}$. Find $n$. 4 291
+Geometry We have triangle $\triangle ABC$ where $AB=AC$ and $AD$ is an altitude.
+Meanwhile, $E$ is a point on $AC$ such that $AB\parallel DE.$ If $BC=12$ and the area of $\triangle ABC$ is $180,$ what is the area of $ABDE$?
+135 235
+Intermediate Algebra Find the sum of all complex roots of the equation $\frac{1}{x-1}+\frac{1}{x-5}+\frac{1}{x-10}+\frac{1}{x-25}=2,$ given that there are no repeated roots. 43 443
+Number Theory Kirsty needs to hire a plumber to fix her house.
+The plumber charges $242_{5}$ dollars for every hour of labor and $367_{8}$ dollars for equipment.
+If the plumber works for $3.5_{10}$ hours, how many dollars (in base ten) will Kirsty owe the plumber?
+499 474
+Prealgebra John and Gary are playing a game.
+John spins a spinner numbered with integers from 1 to 20.
+Gary then writes a list of all of the positive factors of the number spun except for the number itself.
+Gary then creates a new spinner with all of the numbers on his list.
+John then spins this spinner, and the process continues.
+The game is over when the spinner has no numbers on it.
+If John spins a 20 on his first spin, what is the maximum number of total spins (including the one he already made) that John can make before the game is over?
+4 618
+Precalculus Let $\mathbf{a}$ and $\mathbf{b}$ be vectors such that $\mathbf{v}=\operatorname{proj}_{\mathbf{a}}\mathbf{v}+\operatorname{proj}_{\mathbf{b}}\mathbf{v}$ for all vectors $\mathbf{v}.$ Enter all possible values of $\mathbf{a}\cdot\mathbf{b},$ separated by commas.
+0 133
+Table 11: Examples of MATH dataset. We give the format of our processed MATH dataset. Meanwhile, the dataset or its URL will be available on our codebase page upon release.

### v3

@@
 However, placing LLMs into specific roles may reduce their reasoning diversity and performance on a few tasks where role dependence is low.
-To alleviate the excessive immersion of the LLM into a specific role, we propose Nash CoT by constructing a competitive system on each path that balances the generation from role-specific LLMs’ and the general LLMs’ generation, thereby ensuring both effective role adoption and diversity in LLM generation further maintaining the performance of multi-path inference while reducing the requirement of the number of inference paths.
+To alleviate the excessive immersion of the LLM into a specific role, we propose Nash CoT by constructing a game system on each path that balances the generation from role-specific LLMs’ and the general LLMs’ generation, thereby ensuring both effective role adoption and diversity in LLM generation further maintaining the performance of multi-path inference while reducing the requirement of the number of inference paths.
 We evaluate Nash CoT across various inference tasks, including Arabic Reasoning, Commonsense Question Answering, and Symbolic Inference, achieving results that are comparable to or better than those of multi-path CoT with the equal number of inference paths.
@@
 Step 2 utilizes the role-immersed LLM and LLM under normal conditions to collect model predictions separately.
-Step 3 filters the responses based on Nash Equilibrium.
+Step 3 filters the responses to ensure the existence of a unique Nash Equilibrium (NE).
 Subsequently, among various CoT-related improvements, modified by utilizing multi-path inference which is represented by Wang et al.
@@
 This can mitigate the constraints imposed by excessive role casting on the reasoning performance, ultimately maintaining the role-immersed multi-path CoT’s reasoning performance in these questions.
-Driven by this motivation, we construct a gaming system on each path, where the LLM in the role-playing context and the LLM in its normal state act as two distinct players.
-When this gaming system formed by these two players reaches a Nash Equilibrium, the reasoning under that path can balance the preferences of both role-playing with that of the LLM’s normal state.
-Furthermore, we can choose the response with the most Nash Equilibrium hits through voting to ensure the performance of LLM reasoning.
+Driven by this motivation, we construct a game system on each path, where the LLM in the role-playing context and the LLM in its normal state act as two distinct players.
+When this game system formed by these two players reaches a unique Nash Equilibrium (NE), the reasoning under that path can balance the preferences of both role-playing with that of the LLM’s normal state.
+Furthermore, we can choose the response with the most NE hits through voting to ensure the performance of LLM reasoning.
 Figure 2: General Performance Comparison. We compare the average performance of, zero-shot, and zero-shot CoT self-consistency (20 Paths) with our Nash CoT (10 Paths) on Mistral-Instruct and GLM4.
@@
 To summarize, our contribution can be summarized as follows:
-  • To our knowledge, we are the first to introduce the concept of Nash Equilibrium into multi-path CoT, with the objective of balancing text generation guided by the role-immersed LLM with that produced by the LLM’s default state.
+  • To our knowledge, we are the first to introduce the concept of NE into multi-path CoT, with the objective of balancing text generation guided by the role-immersed LLM with that produced by the LLM’s default state.
 This innovation enables us to minimize the number of paths required to achieve good results in multi-path inference, while simultaneously preserving the multi-path CoT’s inherent performance advantages.
@@
 Additionally, Munos et al.
-(2023) proposes combining DPO with Nash Equilibrium to ensure the convergence of the last iterated policy.
-Despite our study also utilizing the concept of Nash Equilibrium in the Preference model, the main difference is that we utilize Nash Equilibrium as a standard to pick up the answer instead of optimizing the LLM’s parameters.
+(2023) proposes combining DPO with NE to ensure the convergence of the last iterated policy.
+Despite our study also utilizing the concept of NE in the Preference model, the main difference is that we utilize NE as a standard to pick up the answer instead of optimizing the LLM’s parameters.
 3 Preference Equilibrium in mini-batch inference
@@
 However, multi-path inference lacks a theoretical foundation for determining the optimal number of inference paths, requiring us to increase the number of paths to ensure performance, and further increase the computational resource consumption.
-To reduce the number of paths required for multi-path inference, we utilize the concept of Nash Equilibrium to locally construct a binary gaming system in multi-path inference.
-Specifically, the preference of each valid inference path of the LLM needs to achieve Nash equilibrium with the preferences of the generation guided by the role-immersed LLM.
+To reduce the number of paths required for multi-path inference, we utilize the concept of NE to locally construct a binary game system in multi-path inference.
+Specifically, the preference of each valid inference path of the LLM needs to achieve NE with the preferences of the generation guided by the role-immersed LLM.
 This approach increases the possibility of each path correctly answering the question while maintaining a certain level of robustness (diversity), thereby reducing the number of inference paths required by self-consistency.
@@
 You enjoy approaching problems from a philosophical perspective and arriving at conclusions that align with your values.
-Geographer: You are a geographer with a deep understanding of geographical knowledge. Please approach the given problem from the perspective of a geographer.
 $\cdots$ (other cases have been appended to the Appendix.)
@@
 $\begin{split}\mathcal{P}(\pi_{2}\prec\pi_{1}):=\mathbb{E}_{{}^{y_{1}\sim\pi_{1}(\cdot|x)}_{y_{2}\sim\pi_{2}(\cdot|x)}}\bigg{[}\mathcal{P}(y_{2}\prec y_{1}|x)\bigg{]},\end{split}$ (2)
-subsequently, we imply the necessity and existence of Nash Equilibrium in the Preference model (shown in Equation 1).
+subsequently, we imply the necessity and existence of a unique NE in the Preference model.
 Preference Equilibrium.
-We first discuss why we should construct a gaming system: To increase the accuracy of each inference path, we can select the most suitable template (example 3) for the given question, immersing the LLM into a template-related role.
+We first discuss why we should construct a game system: To increase the accuracy of each inference path, we can select the most suitable template (example 3) for the given question, immersing the LLM into a template-related role.
 This approach helps LLMs to solve problems more effectively, thereby reducing the number of inference paths needed.
 However, this may lead to some issues: If the LLM is excessively immersed in a specific role, it may reduce the diversity of LLM’s generation, and further reduce LLM’s accuracy in solving tasks that have low role dependence.
-To address these problems, we locally build a bi-player gaming system that the preference of role-immersed LLM (player 1) over the normal status of LLM (player 2) is the pay-off of role-immersed LLM, vice visa.
-If player 1 and player 2 reach Nash Equilibrium  [^2] Preference Equilibrium leverages the concepts of Nash Equilibrium.
-Nash equilibrium is proposed by John Nash.
-It is a concept solution where, assuming each participant knows the equilibrium strategies of the other participants, no participant can benefit by changing their own strategy., then the generation can balance the preference of both player 1 and player 2.
+To address these problems, we locally build a bi-player game system that the preference of role-immersed LLM (player 1) over the normal status of LLM (player 2) is the pay-off of role-immersed LLM, vice visa.
+If player 1 and player 2 reach NE, then the generation can balance the preference of both player 1 and player 2.
+Regularized form of preference model.
+[…]
+[…]
+Unlike the trainable RLHF framework proposed by Munos et al., which aims to regularize the training policy to be close to a known reference (safe) policy $\mu$, our inference system is not trainable.
+Instead, the incorporation of this regularization term serves a different purpose: Facilitate a theoretical analysis that demonstrates that, under the KL constraint, the game system possesses a unique Nash Equilibrium (NE).
 Definition 3.1 (Preference Equilibrium).
-Given any two policys (players) $\pi_{1}$ and $\pi_{2}$ along with the pay-off of $\pi_{1}$ and $\pi_{2}$ as $\mathcal{P}(\pi_{1}\prec\pi_{2})$ and $\mathcal{P}(\pi_{2}\prec\pi_{1})$, we define the Nash Equilibrium of $\pi_{1}$ and $\pi_{2}$ as Preference Equilibrium.
-Subsequently, we define the status that player 1 and player 2 reach Nash Equilibrium as Preference Equilibrium (Definition 3.1).
-Meanwhile, in Theorem 3.1, we prove the existence of Nash Equilibrium in this system.
-In particular, the strategy of player 1 equal to player 2 is one solution that this system has reached Preference Equilibrium.
-Theorem 3.1 (Existence of Preference Equilibrium).
-Given any two policy (player) $\pi_{1}$ and $\pi_{2}$ within the gaming system defined in Definition 1, where $\pi\in\Pi$.
-$\pi_{1}\equiv\pi_{2}$ is a solution where the gaming system reaches Nash Equilibrium.
+Given any two policy (players) $\pi_{1},\pi_{2}$ along with the pay-off of $\pi_{1},\pi_{2}$ as $\mathcal{P}(\pi_{2}\prec\pi_{1})-\tau\cdot{\rm KL}(\pi_{1}||\mu)$ and $\mathcal{P}(\pi_{1}\prec\pi_{2})-\tau\cdot{\rm KL}(\pi_{2}||\mu)$, we define the NE of $\pi_{1}$, $\pi_{2}$ as Preference Equilibrium.
+Subsequently, we define the status that player 1 and player 2 reach NE as Preference Equilibrium (Definition 3.1).
+Meanwhile, in Theorem 3.1, we prove the uniqueness of NE in this system.
+Specifically, having the strategy of player 1 equal to player 2 is a crucial prerequisite for ensuring that the game system has a unique Preference Equilibrium.
+Theorem 3.1 (Uniqueness of Preference Equilibrium).
+Given any two policy (player) $\pi_{1}$ and $\pi_{2}$ within the game system defined in Definition 1, where $\pi\in\Pi$.
+$\pi_{1}=\pi_{2}$ is an essential condition to guarantee this system has a unique NE.
 Proof of Theorem 3.1 see Appendix F.
 Meanings of the existence of Preference Equilibrium.
-Theorem 3.1 proves the existence of a Nash Equilibrium between the role-immersed LLM and the normal status of LLM generation.
-Therefore, when reaching Preference Equilibrium, the preference of generations made by the role-immersed LLM is aligned with those made by the LLM without templates.
-Meanwhile, the preference for role-immersed LLM generation is much closer to the selected role-immersed context, while those of the normal status LLM are predominantly based on its parameters which are much more diverse than role-immersed LLM generation.
+When reaching Preference Equilibrium, the preference of generations made by the role-immersed LLM is aligned with those made by the LLM without templates.
+Specifically, the preference for role-immersed LLM generation is much closer to the selected role-immersed context, while those of the normal status LLM are predominantly based on its parameters which are much more diverse than role-immersed LLM generation.
 Therefore, Preference Equilibrium can balance the requirement of role-immersed LLM generation and the diversity of the LLM generation in each path.
-In particular, $\pi_{1}\equiv\pi_{2}$ means their outputs are also likely to be equal.
-This insight forms a fundamental basis for practically implementing Nash CoT.
+Meanings of $\pi_{1}=\pi_{2}$.
+Theorem 3.1 proves $\pi_{1}=\pi_{2}$ is an essential condition that guaranteeing this game system has a unique NE.
+Therefore, we can choose $\pi_{1}=\pi_{2}$ as the indicator to ensure that this game system has a unique NE.
+Meanings of KL terms.
+If the payoff of $\pi$ lacks a KL term, proving the uniqueness of the Nash Equilibrium (NE) becomes infeasible.
+Furthermore, the KL term plays a crucial preferential role in penalizing role-immersed generation, particularly when it diverges significantly from normal generation.
+This penalty helps to strike a balance between role-immersed and normal generations (see the Proof of Theorem 3.1).
 3.1 Mini-batch inference with Preference Equilibrium
-Subsequently, based on Preference Equilibrium, we conceptualize a mini-batch inference (shown in Step 2 of Figure 1) as a bi-player gaming system.
+Subsequently, based on Preference Equilibrium, we conceptualize a mini-batch inference (shown in Step 2 of Figure 1) as a bi-player game system.
 This approach aims to maintain the performance benefits of multi-path inference while retaining some of the inherent randomness (diversity) of standard inference methods.
@@
 […]
-Based on the mini-batch inference, we further introduce Nash CoT in the next chapter.
-(In particular, in the current framework, successful inference cannot always be guaranteed.
-For instance, $y^{*}$ may not always in $[y_{1},y_{2}]$.
-We will address this problem in the following chapters.)
+[…]
 4 Nash Chain of Thought (Nash CoT)
@@
 This leads us to another type of loop: Outer Loops ($\rm n_{\rm outer}$): This loop resembles the concept of multi-path inference that iterates $\rm n_{\rm mini}$ multiple times.
-After completing loop $\rm n_{\rm outer}$, we filter the generated answers and retain the answer that reaches Preference Equilibrium most frequently (shown in Algorithm 2), as the predicted answer.
-Core LLM Methods SingleEQ AddSub MultiArith GSM8K AQuA SVAMP Avg.
-Mistral-Instruct (7B) zero-shot 15.3$\pm$0.8 12.0$\pm$2.8 3.3$\pm$1.3 2.7$\pm$2.0 20.8$\pm$1.5 7.7$\pm$2.0 10.3$\pm$7
-zero-shot CoT 76.0$\pm$0.8 82.5$\pm$2.0 75.4$\pm$6.1 44.3$\pm$ 4.0 27.9$\pm$2.3 63.4$\pm$6.9 61.6$\pm$19
-self-consistency (20 Paths) 82.5$\pm$0.8 86.3$\pm$5.1 86.3$\pm$2.8 58.5$\pm$2.8 34.4$\pm$6.1 76.5$\pm$2.8 70.8$\pm$19
-Nash CoT (10 Paths) 81.4$\pm$0.8 86.3$\pm$6.0 86.3$\pm$4.7 55.7$\pm$5.8 39.9$\pm$5.4 77.0$\pm$3.5 71.1$\pm$17
-GLM4-chat (9B) zero-shot 1.1$\pm$1.5 1.1$\pm$1.5 12.6$\pm$3.9 12.0$\pm$2.0 22.4$\pm$4.1 4.4$\pm$2.8 8.9$\pm$8
-zero-shot CoT 90.7$\pm$1.5 90.7$\pm$1.5 98.4$\pm$1.3 80.9$\pm$2.8 20.8$\pm$ 3.1 86.9$\pm$3.5 78.1$\pm$26
-self-consistency (20 Paths) 92.3$\pm$2.0 90.2$\pm$2.3 98.4$\pm$2.3 89.3$\pm$0.2 20.8$\pm$3.1 91.5$\pm$1.1 80.4$\pm$27
-Nash CoT (10 Paths) 91.3$\pm$0.8 90.2$\pm$2.7 96.7$\pm$3.3 80.3$\pm$1.3 20.8$\pm$3.1 88.0$\pm$2.0 77.9$\pm$26
-Table 1: Experimental results on Arabic Reasoning benchmarks.
-We test Zero-Shot CoT and Nash CoT with the core LLM includes Mistral-Instruct (7B) and GLM4-chat (9B) on mathematical benchmarks including AddSub, MultiArith, SingleEQ, SVAMP, GSM8K, and AQuA.
-Nash CoT performs the best.
+After completing loop $\rm n_{\rm outer}$, We filter the generated answers and retain the one that satisfies the uniqueness of the NE most frequently.
+(shown in Algorithm 2), as the predicted answer.
 Algorithm 2  Nash CoT (Answer Filtering)
@@
 12:  end if
+Core LLM Methods SingleEQ AddSub MultiArith GSM8K AQuA SVAMP Avg.
+Mistral-Instruct (7B) zero-shot 15.3$\pm$0.8 12.0$\pm$2.8 3.3$\pm$1.3 2.7$\pm$2.0 20.8$\pm$1.5 7.7$\pm$2.0 10.3$\pm$7
+zero-shot CoT 76.0$\pm$0.8 82.5$\pm$2.0 75.4$\pm$6.1 44.3$\pm$ 4.0 27.9$\pm$2.3 63.4$\pm$6.9 61.6$\pm$19
+self-consistency (20 Paths) 82.5$\pm$0.8 86.3$\pm$5.1 86.3$\pm$2.8 58.5$\pm$2.8 34.4$\pm$6.1 76.5$\pm$2.8 70.8$\pm$19
+Nash CoT (10 Paths) 81.4$\pm$0.8 86.3$\pm$6.0 86.3$\pm$4.7 55.7$\pm$5.8 39.9$\pm$5.4 77.0$\pm$3.5 71.1$\pm$17
+GLM4-chat (9B) zero-shot 1.1$\pm$1.5 1.1$\pm$1.5 12.6$\pm$3.9 12.0$\pm$2.0 22.4$\pm$4.1 4.4$\pm$2.8 8.9$\pm$8
+zero-shot CoT 90.7$\pm$1.5 90.7$\pm$1.5 98.4$\pm$1.3 80.9$\pm$2.8 20.8$\pm$ 3.1 86.9$\pm$3.5 78.1$\pm$26
+self-consistency (20 Paths) 92.3$\pm$2.0 90.2$\pm$2.3 98.4$\pm$2.3 89.3$\pm$0.2 20.8$\pm$3.1 91.5$\pm$1.1 80.4$\pm$27
+Nash CoT (10 Paths) 91.3$\pm$0.8 90.2$\pm$2.7 96.7$\pm$3.3 80.3$\pm$1.3 20.8$\pm$3.1 88.0$\pm$2.0 77.9$\pm$26
+Table 1: Experimental results on Arabic Reasoning benchmarks.
+We test Zero-Shot CoT and Nash CoT with the core LLM includes Mistral-Instruct (7B) and GLM4-chat (9B) on mathematical benchmarks including AddSub, MultiArith, SingleEQ, SVAMP, GSM8K, and AQuA.
+Nash CoT performs the best.
 Answer Filtering.
-In terms of answer filtering, as shown in Algorithm 2 we first count the most frequent prediction to satisfy Preference equilibrium.
+In terms of answer filtering, as shown in Algorithm 2 we first count the most frequent prediction to satisfy the uniqueness of the NE.
 Specifically, we count all $y^{*}$ satisfy $y^{*}\in[y_{1},y_{2}]$ and compute their frequency.
@@
 (2022).
-In particular, all of these selected LLMs are turned via RLHF, and the difference between LLM turned with RLHF and the original foundation models have been detailed by Ouyang et al..
+In particular, all of these selected LLMs are turned via RLHF, and the difference between LLM turned with RLHF and the original foundation models have been detailed by Ouyang et al.
 Baselines.
@@
 Additionally, we have provided the Player Templates $x^{t}$ in Table 1 and the Appendix, meanwhile, we utilizing the same CoT template $x^{c}$ as in zero-shot CoT.
-Specifically, in Table 12 and 3, we set the total number of reasoning paths for Nash CoT to 10 and for self-consistency to 20.
+In particular, in Table 12 and 3, we set the total number of reasoning paths for Nash CoT to 10 and for self-consistency to 20.
 The primary goal is to demonstrate that Nash CoT can achieve the same or even better results with only half the number of reasoning paths compared to self-consistency.
@@
 2) Does the template improve the accuracy of path predictions, and what impact does it have on experimental performance?
-3) Is the Nash Equilibrium the reason for the performance improvement observed in Nash CoT?
+3) Is the NE the reason for the performance improvement observed in Nash CoT?
 As the number of inference paths increases, Nash CoT can obviously surpass self-consistency with fewer inference paths.
@@
 Table 6: Results on GSM8K. The templates for these roles are shown in Table 9 on the Appendix.
-Effectiveness of Nash Equilibrium.
-While Nash CoT has demonstrated effectiveness across various LLMs and benchmarks, however, the extent to which Nash CoT’s performance is driven by Nash Equilibrium remains to be determined.
+Effectiveness of NE.
+While Nash CoT has demonstrated effectiveness across various LLMs and benchmarks, however, the extent to which Nash CoT’s performance is driven by NE remains to be determined.
 To further investigate, we compare the self-consistency integrated template (denoted as self-con.+temp (10 paths)) with Nash CoT (10 paths) using Mistral-Instruct (7B).
@@
 Nash CoT 39.9 77.0 55.7 86.3 64.7
-Table 7: Ablation of Nash Equilibrium. This experiment is conducted to verify the necessity of Nash Equilibrium in the performance of Nash CoT.
+Table 7: Ablation of NE. This experiment is conducted to verify the necessity of NE in Nash CoT.
 7 Conclusion
-We prove the existence of a Nash equilibrium in preferences between role-immersed and normal-status LLM generations, and develop a gaming system within each path of the multi-path CoT to introduce Nash CoT.
+We prove the existence of a NE in preferences between role-immersed and normal-status LLM generations, and develop a game system within each path of the multi-path CoT to introduce Nash CoT.
 Experimental results show that Nash CoT can perform equally or even better than self-consistency while requiring fewer paths.
@@
 Meanwhile, we thank the reviewers for their valuable suggestions.
-Ethics Claims
-Pre-training or fine-tuning an LLM requires much more computing resources.
-Multi-path CoT is an ideal approach that has been proven to enhance the performance of LLMs’ inference.
-Meanwhile, Nash CoT effectively reduces the inference paths needed and times of multi-path inference.
-We believe our approach can further elevate the effectiveness of multi-path inference, thereby further improving the effectiveness of LLM.
 Appendix A Dataset
@@
 Appendix F Proof of Theorem 3.1.
-Subsequently, we prove the existence of Nash equilibrium in this system.
-For any two given polices $\pi_{1}\in\Pi$ and $\pi_{2}\in\Pi$ We first define the pay-off of $\pi_{1}$ and $\pi_{2}$ as $R(\pi_{1};\pi_{2})$ and $R(\pi_{2};\pi_{1})$:
-$\begin{split}R(\pi_{1};\pi_{2})=\mathcal{P}(\pi_{1}\prec\pi_{2})\\ R(\pi_{2};\pi_{1})=\mathcal{P}(\pi_{1}\succ\pi_{2}),\end{split}$ (3)
-we provide the proof of the existence of Nash equilibrium in this system.
-We define $\bar{\pi}=[\pi_{1},\pi_{2}]$, $v(\bar{\pi})=[R(\pi_{1};\pi_{2}),R(\pi_{1};\pi_{2})]$.
-According to the Nash equilibrium, it should have to satisfy this relationship:
-$\begin{split}v(\bar{\pi}^{*})(\bar{\pi}^{*}-\bar{\pi})\leq 0\end{split}$ (4)
-Subsequently, refer to Munos et al., we can learn that if we want Equation 4 holds true, we just have to guarantee Equation 5 holds true.
-$\begin{split}(v(\bar{\pi})-v(\bar{\pi}^{\prime}))^{\rm T}(\bar{\pi}-\bar{\pi}^{\prime})\leq 0,\end{split}$ (5)
-where $\bar{\pi}$ and $\bar{\pi}^{\prime}$ are any two given policy set. Subsequently, we can further arrive at the following relationships:
-$\begin{split}(v(\bar{\pi})-v(\bar{\pi}^{\prime}))^{\rm T}(\bar{\pi}-\bar{\pi}^{\prime})&=\begin{pmatrix}R(\pi_{1};\pi_{2})-R(\pi_{1}^{\prime};\pi_{2}^{\prime})\\ R(\pi_{2};\pi_{1})-R(\pi_{2}^{\prime};\pi_{1}^{\prime})\\ \end{pmatrix}\cdot\begin{pmatrix}\pi_{1}-\pi_{1}^{\prime},\pi_{2}-\pi_{2}^{\prime}\\ \end{pmatrix}\\ &=\bigg{(}R(\pi_{1};\pi_{2})-R(\pi_{1}^{\prime};\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime})+\bigg{(}R(\pi_{2};\pi_{1})-R(\pi_{2}^{\prime};\pi_{1}^{\prime})\bigg{)}\\ &\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime})+\bigg{\{}2-\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})+\\ &\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\bigg{\}}\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime}-\pi_{2}+\pi_{2}^{\prime})+2\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})-2\bigg{)}\cdot(\pi_{2}^{\prime}-\pi_{2})+\\ &\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime}).\end{split}$ (6)
-In particular, we can find that if $\bar{\pi}\equiv\bar{\pi}^{\prime}$ then $(v(\bar{\pi})-v(\bar{\pi}^{\prime}))^{\rm T}(\bar{\pi}-\bar{\pi}^{\prime})\equiv 0$, thus $\bar{\pi}\equiv\bar{\pi}^{\prime}$ is one solution that $\pi_{1}$ and $\pi_{2}$ has reached equilibrium.
-Appendix G Further clarify our settings
+Our proof is refereed to Munos et al.
+Here, we examine a preference model incorporating a Kullback-Leibler (KL) divergence constraint.
+Unlike the trainable system proposed by Munos et al., which aims to serve as the policy generating the training data for the preference model or address scenarios requiring adherence to a known safe policy, our inference system is non-trainable.
+Consequently, a static distance is maintained between the players $(\pi_{1},\pi_{2})$ and a static reference policy $\mu$.
+The introduction of this regularization term is intended to facilitate a theoretical analysis demonstrating that, under the KL constraint, the game system possesses a unique Nash Equilibrium (NE).
+[…]
+$\begin{split}\mathcal{P}_{\tau}(\pi_{1}\prec\pi_{2})\overset{def}{=}\mathcal{P}(\pi_{1}\prec\pi_{2})-\tau\cdot{\rm KL}(\pi_{2}(\cdot|x)||\mu(\cdot|x))+\tau\cdot{\rm KL}(\pi_{1}(\cdot|x)||\mu(\cdot|x)).\end{split}$ (3)
+Then we prove the existence of NE.
+Specifically, the mappings $\pi_{1}\mapsto\mathcal{P}(\pi_{1}\prec\pi_{2})$ and $\pi_{2}\mapsto\mathcal{P}(\pi_{1}\prec\pi_{2})$ are linear in $\pi_{1}$ (respectively in $\pi_{2}$) thus $\pi_{1}\mapsto\mathcal{P}_{\tau}(\pi_{1}\prec\pi_{2})$ is concave and $\pi_{2}\mapsto\mathcal{P}_{\tau}(\pi_{1}\prec\pi_{2})$ is convex.
+Existence of a NE is derived from the mini-max theorem for convex-concave functions Sion (1958).
+Next, we substitute Equation 3 into our problem setup.
+In our setup, model training is not involved, so $\mu$ can be regarded as one of the strategies in $\pi_{1}$ and $\pi_{2}$.
+Furthermore, we assume $\mu$ to be $\pi_{1}$, and $\pi_{2}$ is the role immersed policy, so we can rewrite the Equation 3 as a two-player game system, with the pay-off of $\pi_{1}$ and $\pi_{2}$:
+$\begin{split}R(\pi_{1};\pi_{2})&=\mathcal{P}(\pi_{2}\prec\pi_{1})-\tau\cdot{\rm KL}(\pi_{1}||\mu)\\ &=\mathcal{P}(\pi_{2}\prec\pi_{1})\end{split}$ (4)
+and
+$\begin{split}R(\pi_{2};\pi_{1})&=\mathcal{P}(\pi_{1}\prec\pi_{2})-\tau\cdot{\rm KL}(\pi_{2}||\mu)\\ &\leq\mathcal{P}(\pi_{1}\prec\pi_{2})\end{split}$ (5)
+In particular, when the difference between $\pi_{2}$ and $\mu$ is small enough, R($\pi_{2}$; $\pi_{1}$) can be approximately regarded as $\mathcal{P}(\pi_{1}\prec\pi_{2})$.
+When the difference between $\pi_{2}$ and $\mu$ is large enough, $\mathcal{P}(\pi_{1}\prec\pi_{2})$ serves as an upper bound for $R(\pi_{2};\pi_{1})$.
+Therefore, the KL term plays a constraining term in the role-immersed generation process.
+Subsequently, to show this game has a unique NE, we have to show the variational inequality is strictly monotone (Rosen, Theorem 2).
+Meanwhile, followed by the proof of Munos et al.
+We learn that every NE of the game should satisfy the variational inequality for all $\bar{\pi}$:
+$\begin{split}v^{T}(\bar{\pi}^{*})(\bar{\pi}^{*}-\bar{\pi})\leq 0,\end{split}$ (6)
+Furthermore, the inequality is strictly monotone if and only if for every $\bar{\pi}_{1}$ and $\bar{\pi}_{2}$ we have that:
+$\begin{split}(v(\bar{\pi}_{1})-v(\bar{\pi}_{2}))^{T}(\bar{\pi}_{1}-\bar{\pi}_{2})\leq 0,\end{split}$ (7)
+And this inequality has been proved by Munos et al., with equality only at $\bar{\pi}_{1}=\bar{\pi}_{2}$.
+Therefore, $\pi_{1}=\pi_{2}$ is an essential condition for maintaining strictly monotone.
+Meanwhile, strictly monotone serves as the guarantee of the uniqueness of the NE.
+Therefore, we selected $\pi_{1}=\pi_{2}$ as an indicator to ensure the uniqueness of the NE.
+Appendix G Extra Explanation
 Why in Table 12 and 3 do we set the number of pathways for Nash CoT to be half of self-consistency?